### v3

# Common Task Framework For a Critical Evaluation of Scientific Machine Learning Algorithms

Abstract
Machine learning (ML) is transforming modeling and control in the physical, engineering, and biological sciences. However, rapid development has outpaced the creation of standardized, objective benchmarks—leading to weak baselines, reporting bias, and inconsistent evaluations across methods. This undermines reproducibility, misguides resource allocation, and obscures scientific progress. To address this, we develop a Common Task Framework (CTF) for scientific machine learning. The CTF features a curated set of datasets and task-specific metrics spanning forecasting, state reconstruction, and generalization under realistic constraints, including noise and limited data. Inspired by the success of CTFs in fields like natural language processing and computer vision, our framework provides a structured, rigorous foundation for head-to-head evaluation of diverse algorithms. As a first step, we benchmark methods on two canonical nonlinear systems: Kuramoto-Sivashinsky and Lorenz. These results illustrate the utility of the CTF in revealing method strengths, limitations, and suitability for specific classes of problems and diverse objectives. Next, we are launching a competition based on a global, real-world sea surface temperature dataset with a true holdout dataset to foster community engagement. Our long-term vision is to replace ad hoc comparisons with standardized evaluations on hidden test sets, thereby raising the bar for rigor and reproducibility in scientific ML.
1 Introduction
Data science, especially machine learning (ML) and artificial intelligence (AI), is transforming almost every aspect of the engineering, physical, social, and biological sciences. As the body of literature on new ways to model many scientific data and systems grows, we still lack objective measures to adequately characterize and compare these methods. In the absence of a common standard for benchmarking new and existing approaches, the current literature suffers from weak baselines, reporting bias, and inconsistent evaluations McGreivy2024, thus leading to a clear call for a CTF for evaluating machine learning methods on a diverse set of metrics related to science and engineering kutz2025accelerating. Several benchmark frameworks have been proposed to address this gap in scientific machine learning. For example, The Well ohana2024well provides a large-scale collection of diverse physics simulation datasets across multiple domains. CoDBench burark2024codbench offers a comprehensive benchmarking suite to systematically evaluate data-driven models for solving differential equations and continuous dynamical systems. PDEBench takamoto2022pdebench and PDEArena gupta2022towards are PDE-focused benchmarking frameworks that provide curated datasets and task suites to assess the accuracy and efficiency of ML-based solvers. These benchmarks exemplify the move toward standardized, reproducible evaluation in scientific ML. Nevertheless, despite the rise of benchmark data sets across science and engineering, the reliance on self-reporting has generated a significant reproducibility crisis. Self-reporting is, in general, a flawed premise. For instance, neural networks upon training are typically initialized with random weight assignment. Although the errors achieved on the training data set are comparable from run to run, the errors on the test set can be significantly different. This can lead to $p$-hacking, or judicious picking of results, when reporting scores on test data sets, i.e. simply re-train the model until a desired and good result is achieved for self-reporting. Only with a true, withheld test set is a comparison among methods possible.
CTFs play a critical role in evaluating methodological advancements. Donoho donoho2017 has argued that the successful application of CTFs is a primary factor for the success of data science and machine learning. Indeed, the fields of speech recognition, natural language processing, and computer vision have developed mature CTF platforms that are progressively updated with more challenging data in order to drive progress and innovation. For instance, the industry-leading CVPR conference offers more than 30 challenge problems per year for participants to score and benchmark their ML/AI algorithms against. More broadly, classic fields of machine learning have benefited from extensive benchmark environments and common task frameworks, including ImageNet deng2009imagenet; krizhevsky2012imagenet, Go and chess silver2018general, video games such as Atari mnih2015human and StarCraft vinyals2019grandmaster, the OpenAI Gym ravichandiran2018hands; dutta2018reinforcement, among other environments for more realistic control deisenroth2011pilco; todorov2012mujoco. Unlike these leading fields, many scientific disciplines have yet to integrate the CTF into their core infrastructure McGreivy2024. This compromises true comparative metrics between methods, algorithms, and results, and it limits the potential of ML in these areas.
Figure 1: The twelve-axis radar plot characterizes a method’s performance across all tasks on a dataset, and provides a visual performance profile. The axes correspond to the various tasks associated with forecasting and reconstruction with noise, limited data and parametric dependency. The chart shows the top four performing metrics on the KS and the Lorenz dataset scored against their reference baselines: constant zero and average prediction respectively.
1.1 Common Task Framework for Science and Enginering
We propose a CTF for science and engineering that is primarily focused on evaluating machine learning and AI models for dynamic systems: systems whose underlying evolution is determined by physical or biophysical principles or governing equations. The CTF will provide training data sets with clear and concise goals related to forecasting and reconstruction under various challenging scenarios, such as noisy measurements, limited data, or varying system parameters, as advocated in kutz2025accelerating. Given a training dataset and a range of timesteps to predict, users will produce approximations for a hidden test dataset. The predictions are evaluated and scored on a diverse set of metrics by an independent referee and posted on a leaderboard.
Scoring is by nature reductive—reducing a method’s performance to a single floating point value. We choose a multi-metric scoring approach because a single number often doesn’t provide enough information on whether a method is right for an application or not. As a result, we decided to carefully design a twelve-score system designed to match crucial tasks required in science and engineering. A summary, or composite score, is also produced that gives the overall score for a given method. Rankings by task and overall performance are highlighted here and tracked on a leader board.
To visualize the overall performance of a method, a radar plot is generated highlighting the various scores associated with the challenge (see Fig.  1). From this figure one can glean the characterization of a method with respect to its performance on the diverse set of CTF tasks. The average of all scores serves as the composite score. This scoring system prevents a winner takes all framework, since different modeling approaches will excel on different tasks. Some will do well with noise, others will not. Others might excel in the limited data regime, while performing poorly under parametric generalization. These profiles are important to provide a comprehensive and well-rounded performance metric, and help guide for scientists for selecting a suitable method.
Once the ctf4science is launched [^1] Kaggle launch date TBD, we invite everyone to benchmark their methods on the CTF for Science by taking the following steps:
  1. 1.
    Sign-up and Sign-in on Kaggle
  2. 2.
    Train your model with our training data and generate predictions for each benchmark case
  3. 3.
    Submit prediction files to the competition platform
  4. 4.
    See your score on the leaderboard
To interact with ctf4science before the competition launch visit our GitHub repository [^2] Available at https://github.com/CTF-for-Science/ctf4science, install the ctf4science packagewydercommon, and evaluate your method on our datasets ODE_Lorenz, PDE_KS, and SST. Our datasets and our ctf4science Python package don’t require high-performance hardware and can be run on a laptop computer.
Figure 2: The CTF Evaluation framework scores the performance of methods on (a) the Lorenz dynamical system and (b) the Kuramoto-Sivashinsky partial differential equation. (c) Data is collected and organized into matrices which is then split into testing and training sets. RMSE errors are computed for reconstruction and short-time forecasting, while the spectral error computes the statistics of long-time forecasting (spatial or temporal). (d) Forecasting and reconstruction tasks are evaluated on noise-free, low-noise and high-noise data. Methods are also evaluated when (e) only limited data is available and (f) for reconstruction of parametrically dependent data.
2 Datasets & Evaluation Metrics
We launch the CTF platform with two canonical and commonly used models in scientific machine learning: the Lorenz equations, a dynamical system and the Kuramoto-Sivashinsky (KS) equation, a partial differential equation. Both exhibit complex and challenging behavior for the science and engineering tasks of reconstruction and forecasting under the constraints of noise, limited data, and parametric dependence. While these equations serve as a starting point, the CTF will evolve to include both more complex data and more challenging tasks. The CTF framework is a sustainable platform that evolves and grows as the community develops more sophisticated methods and algorithms and faces new challenges.
We provide a detailed breakdown of the evaluation metrics and the associated data matrices in the following sections. For convenience, we included an overview table that summarizes the relationship between each evaluation metric and the corresponding data matrices in the supplementary materials.
2.1 Spatio-Temporal System: Kuramoto-Sivashinsky
The KS equation is a fourth order, nonlinear partial differential equation. It is considered a canonical example of spatio-temporal chaos in a one-dimensional PDE and is therefore commonly used as a test problem for data-driven algorithms. The KS equation is a particularly challenging case for fitting algorithms due to its combination of high dimensionality, nonlinearity, and sensitivity to initial conditions (chaotic behavior):
$u_{t}+uu_{x}+u_{xx}+\mu u_{xxxx}=0.$ (1)
The solutions of Eq. (1) are defined on a grid across the domain of $0,32\pi]$ with periodic boundary conditions. A numerical integrator with an unknown time step $\Delta t$ evolves the solution $m$ steps.
2.1.1 Test 1: Forecasting (2 scores)
The first test of the method, as illustrated in Fig. 2-d, involves the approximation of the future state of the system. Thus, given a data matrix representing the dynamics over $t\in[0,10T]$ (${\bf X}_{1}\in\mathbb{R}^{10m\times n}$), the forecast is requested for $t\in[10T,11T]$ (${\bf X}_{1\text{pred}}\in\mathbb{R}^{m\times n}$), with $n$ being the dimension of the system and $m$ being the number of time steps. The forecasting score is composed of two scores evaluating both the short-time forecast $E_{{\mbox{\tiny ST}}}$ (the "weather forecast"), which is computed using root-mean square error (RMSE) between the test set and the user’s approximation, and the long-term forecast $E_{{\mbox{\tiny LT}}}$ (the "climate forecast"), which is based upon the power spectral density - see Fig. 2-c. As such, the following two error scores are computed:
[…]
For the challenge dynamics of interest, sensitivity of initial conditions is common, making long range forecasting to match the test set an unreasonable task given fundamental mathematical limitations with Lyapunov times. Thus, as shown above, the long-time error is computed by least-squares fitting of the power spectrum ${\bf P}[k,:]=\ln(|\mbox{FFT}({{\bf X}[k,:]})|^{2})$, where the fftshift has been used to model the data in the wavenumber domain and ${\bf k}=n/2-k_{max}:n/2+(k_{max}+1)$ with $k_{max}=100$. This means that we look at the match in the first 100 wavenumbers of the power spectrum over a long time simulation. It is clear that there are many ways to evaluate the long-range forecasting capabilities. We chose a simple and transparent metric fully understanding that more nuanced scoring could be used. To provide a reasonable range we then compute the two scores
$E_{1}=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{1\text{pred}},{\bf X}_{1\text{test}})),\quad E_{2}=100(1-S_{{\mbox{\tiny LT}}}({\bf X}_{1\text{pred}},{\bf X}_{1\text{test}})),$ (4)
meaning in each case a score of $E_{i}=100$ corresponds to a perfect match. Note that, as a baseline, a solution guess of zeros $\tilde{\bf X}_{1\text{pred}}[1:k,:]={\bf 0}$ (corresponding also to $\tilde{\bf P}_{1\text{pred}}[N-k:N,{\bf k}]={\bf 0}$) gives a score of $E_{1}=E_{2}=0$.
Input: ${\bf X}_{1\text{train}}\in\mathbb{R}^{10m\times n}$; Output: ${\bf X}_{1\text{pred}}\in\mathbb{R}^{m\times n}$; Scores: $E_{1},E_{2}$.
2.1.2 Test 2: Noisy Data (4 scores)
The ability to handle noise is critical in all data-driven applications as sensors and measurement technologies are by default embedded with varying levels of noise. Methods that work with numerically accurate data, for example data points that are $10^{-6}$ accurate, may be useful for model reduction, but are rarely suitable for discovery and engineering design from real-world data. Both strong and weak noise are considered as these represent realistic challenges to be addressed in practice.
This test is very similar to Test 1, but now with noise added to the data. Specifically, the challenger is given a data matrix ${\bf X}_{2\text{train}}\in\mathbb{R}^{10m\times n}$ and ${\bf X}_{3\text{train}}\in\mathbb{R}^{10m\times n}$ representing the evolution with medium or high noise respectively. The objective is to first produce a reconstruction of the data itself, i.e. denoise the data to produce an estimate of the true state of the dynamics, ${\bf X}_{2\text{pred}},{\bf X}_{4\text{pred}}\in\mathbb{R}^{10m\times n}$ for ${\bf X}_{2\text{train}},{\bf X}_{3\text{train}}$ respectively, and the second objective is to then forecast the future state, matrices ${\bf X}_{3\text{pred}},{\bf X}_{5\text{pred}}\in\mathbb{R}^{m\times n}$ for ${\bf X}_{2\text{train}},{\bf X}_{3\text{train}}$ respectively. For the first task, a least-square fit is used between the approximation of the denoised data and the truth, and for the forecasting a long-time evaluation is computed leading to the following scores:
$E_{3}$ $=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{2\text{pred}},{\bf X}_{2\text{test}})),\quad E_{4}=100(1-S_{{\mbox{\tiny LT}}}({\bf X}_{3\text{pred}},{\bf X}_{3\text{test}})),$
$E_{5}$ $=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{4\text{pred}},{\bf X}_{4\text{test}})),\quad E_{6}=100(1-S_{{\mbox{\tiny LT}}}({\bf X}_{5\text{pred}},{\bf X}_{5\text{test}})).$
Input: ${\bf X}_{2\text{train}},{\bf X}_{3\text{train}}\in\mathbb{R}^{10m\times n}$;  Output: ${\bf X}_{2\text{pred}},{\bf X}_{4\text{pred}}\in\mathbb{R}^{10m\times n}$, ${\bf X}_{3\text{pred}},{\bf X}_{5\text{pred}}\in\mathbb{R}^{m\times n}$; Scores: $E_{3},E_{4},E_{5},E_{6}$.
2.1.3 Test 3: Limited Data (4 scores)
Data limitations are common in real world physical systems and often affect the success of data-driven methods. Thus, testing for model performance on low-data is critically important and provides important insight to potential users.
Figure 2-e demonstrates the nature of the test. In this case only a limited number of snapshots $M$ on numerically accurate data are given ${\bf X}_{4\text{train}}\in\mathbb{R}^{M\times n}$. From this limited data, a forecast must be made which is evaluated with both error metrics (2) & (3) on the approximated future ${\bf X}_{6\text{pred}}\in\mathbb{R}^{m\times n}$. The experiment is repeated with noise on the measurements using the training matrix ${\bf X}_{5\text{train}}\in\mathbb{R}^{M\times n}$ for which a forecasting prediction matrix is produced ${\bf X}_{7\text{pred}}\in\mathbb{R}^{m\times n}$. The performance is evaluated on the following scores representing short and long-time metrics for both noise-free and noisy data respectively.
$E_{7}$ $=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{6\text{pred}},{\bf X}_{6\text{test}})),\quad E_{8}=100(1-S_{{\mbox{\tiny LT}}}({\bf X}_{6\text{pred}},{\bf X}_{6\text{test}})),$
$E_{9}$ $=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{7\text{pred}},{\bf X}_{7\text{test}})),\quad E_{10}=100(1-S_{{\mbox{\tiny LT}}}({\bf X}_{7\text{pred}},{\bf X}_{7\text{test}})).$
Two error scores (analogous to $E_{1}$ and $E_{2}$) are produced for the noise-free and noisy limited data. These scores are $E_{7}$ (short) and $E_{8}$ (long) for the noise free case and $E_{9}$ (short) and $E_{10}$ (long) for the noisy case.
Input: ${\bf X}_{4\text{train}},{\bf X}_{5\text{train}}\in\mathbb{R}^{M\times n}$; Output: ${\bf X}_{6\text{pred}},{\bf X}_{7\text{pred}}\in\mathbb{R}^{m\times n};$Scores: $E_{7},E_{8},E_{9},E_{10}$.
2.1.4 Test 4: Parametric Generalization (2 scores)
Finally, the ability of a model to generalize to different parameter values is evaluated. For this case, the model’s ability to interpolate and extrapolate to new parameter regimes is considered with noise-free data and noisy data respectively. The interpolation and extrapolation are each their own score. This gives a total of four scores that evaluate parametric dependence.
Figure 2-f shows the basic architecture of the test. For the noise-free case, three training data sets are provided with three different (unknown) parameter values ${\bf X}_{6\text{train}},{\bf X}_{7\text{train}},{\bf X}_{8\text{train}}\in\mathbb{R}^{10m\times n}$. Construction of the dynamics in parametric regimes that are interpolatory ${\bf X}_{8\text{pred}}\in\mathbb{R}^{m\times n}$ and extrapolatory ${\bf X}_{9\text{pred}}\in\mathbb{R}^{m\times n}$ are required. For both of the test regimes, a burn in matrix ${\bf X}_{9\text{train}}$ and ${\bf X}_{10\text{train}}$ respectively of size $M\times n$ is given and the performance is evaluated using the short term metric (2).
$E_{11}$ $=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{8\text{pred}},{\bf X}_{8\text{test}})),\quad E_{12}=100(1-S_{{\mbox{\tiny ST}}}({\bf X}_{9\text{pred}},{\bf X}_{9\text{test}})).$
Input: ${\bf X}_{6\text{train}},{\bf X}_{7\text{train}},{\bf X}_{8\text{train}}\in\mathbb{R}^{10m\times n},{\bf X}_{9\text{train}},{\bf X}_{10\text{train}}\in\mathbb{R}^{M\times n}$};
Output: ${\bf X}_{8\text{pred}},{\bf X}_{9\text{pred}}\in\mathbb{R}^{m\times n}$; Scores: $E_{11},E_{12}$.
Figure 3: Ranked average scores of each model on the KS and Lorenz Dataset.
2.2 Dynamical System: Lorenz
One of the most influential dynamical systems in history, the Lorenz dynamical system is given by
$\frac{dx}{dt}=\sigma(y-x),\quad\frac{dy}{dt}=rx-xz-y,\quad\frac{dz}{dt}=xy-bz.$
where the parameters $b=8/3$ and $\sigma=10$ are typically fixed at these values while $r$ is explored as a bifurcation parameter. For specific values of $r$, including our choice $r=28$, the system exhibits chaotic behavior as shown in Fig. 2(a).
Table 1: Model performances for each metric on each dataset (mean $\pm$ std).
Model Avg Score E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12
LSTM hochreiter1997lstm 64.54 ($\pm$ 0.00) 99.34 ($\pm$ 0.18) 52.37 ($\pm$ 12.34) 97.42 ($\pm$ 0.10) 50.75 ($\pm$ 28.41) 96.44 ($\pm$ 0.10) 72.13 ($\pm$ 3.54) 66.61 ($\pm$ 6.43) 32.75 ($\pm$ 16.91) 36.39 ($\pm$ 0.00) 29.57 ($\pm$ 14.47) 80.58 ($\pm$ 0.00) 60.11 ($\pm$ 0.00)
DeepONet deeponet 57.80 ($\pm$ 0.00) 99.11 ($\pm$ 0.27) 45.28 ($\pm$ 40.54) 96.23 ($\pm$ 0.00) 62.32 ($\pm$ 5.49) 91.84 ($\pm$ 0.71) 14.35 ($\pm$ 43.11) 21.84 ($\pm$ 5.71) 40.75 ($\pm$ 18.96) 34.53 ($\pm$ 2.52) 25.15 ($\pm$ 10.17) 85.52 ($\pm$ 5.57) 76.68 ($\pm$ 13.97)
Reservoir jaeger_echo_no_date; maass_computational_2004; pathak_model-free_2018 54.87 ($\pm$ 0.00) 99.91 ($\pm$ 0.06) 56.72 ($\pm$ 10.00) 97.41 ($\pm$ 0.01) -33.49 ($\pm$ 70.39) 95.07 ($\pm$ 0.02) 18.99 ($\pm$ 47.02) 65.21 ($\pm$ 0.00) 61.92 ($\pm$ 9.21) 45.41 ($\pm$ 0.84) -48.51 ($\pm$ 24.31) 99.80 ($\pm$ 0.12) 99.97 ($\pm$ 0.00)
ODE-LSTM coelho2024odelstm 41.67 ($\pm$ 0.00) 97.77 ($\pm$ 0.95) 17.07 ($\pm$ 48.03) 97.90 ($\pm$ 0.14) 69.92 ($\pm$ 2.74) 96.48 ($\pm$ 0.08) -82.40 ($\pm$ 0.00) 55.82 ($\pm$ 8.48) 15.20 ($\pm$ 28.53) 36.83 ($\pm$ 2.90) 14.56 ($\pm$ 23.22) 39.90 ($\pm$ 0.00) 40.95 ($\pm$ 0.00)
Spacetime zhang2023spacetime 38.79 ($\pm$ 0.00) 19.26 ($\pm$ 0.00) 83.52 ($\pm$ 15.01) -100.00 ($\pm$ 0.00) 12.16 ($\pm$ 66.22) 39.32 ($\pm$ 0.00) 56.00 ($\pm$ 8.78) 28.28 ($\pm$ 0.00) 23.68 ($\pm$ 34.66) 33.12 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 77.32 ($\pm$ 0.00) 54.09 ($\pm$ 0.00)
KAN liu2025kankolmogorovarnoldnetworks 35.32 ($\pm$ 0.00) 82.89 ($\pm$ 26.43) 20.53 ($\pm$ 58.61) 96.19 ($\pm$ 0.06) 55.20 ($\pm$ 3.28) 93.00 ($\pm$ 0.01) -76.80 ($\pm$ 32.81) 50.52 ($\pm$ 8.64) -2.45 ($\pm$ 34.69) 33.68 ($\pm$ 4.41) -31.47 ($\pm$ 38.67) 41.69 ($\pm$ 3.90) 60.85 ($\pm$ 0.17)
SINDy sindy; ensemblesindy 33.08 ($\pm$ 0.00) 81.83 ($\pm$ 0.00) 36.00 ($\pm$ 0.00) 36.85 ($\pm$ 0.00) 69.92 ($\pm$ 2.74) -29.22 ($\pm$ 0.00) -18.27 ($\pm$ 0.00) 55.82 ($\pm$ 8.48) 15.20 ($\pm$ 28.53) 36.83 ($\pm$ 2.90) 14.56 ($\pm$ 23.22) 82.38 ($\pm$ 0.00) 15.07 ($\pm$ 0.00)
OptDMD askham2018variable 20.62 ($\pm$ 0.00) 52.08 ($\pm$ 0.00) -67.33 ($\pm$ 0.00) 55.51 ($\pm$ 0.00) 1.33 ($\pm$ 0.00) 56.85 ($\pm$ 0.00) -64.13 ($\pm$ 0.00) 59.11 ($\pm$ 0.00) 8.67 ($\pm$ 0.00) 42.15 ($\pm$ 0.00) -15.73 ($\pm$ 0.00) 59.23 ($\pm$ 0.00) 59.68 ($\pm$ 0.00)
FNO li2021fourier 20.51 ($\pm$ 0.00) 50.88 ($\pm$ 12.55) -33.65 ($\pm$ 75.34) 54.69 ($\pm$ 0.00) -9.47 ($\pm$ 63.22) 56.40 ($\pm$ 0.00) 35.36 ($\pm$ 35.71) 20.82 ($\pm$ 16.51) 51.36 ($\pm$ 6.19) -100.00 ($\pm$ 100.00) 32.53 ($\pm$ 19.61) 29.29 ($\pm$ 8.73) 57.95 ($\pm$ 11.03)
PyKoopman bruntonkutzkoopmanreview22; Pan2024 8.33 ($\pm$ 0.00) 34.50 ($\pm$ 0.00) 89.87 ($\pm$ 0.00) 54.97 ($\pm$ 0.01) 52.40 ($\pm$ 0.00) 56.48 ($\pm$ 0.00) -90.59 ($\pm$ 0.72) -22.31 ($\pm$ 0.00) -93.73 ($\pm$ 0.00) 43.93 ($\pm$ 0.98) -78.67 ($\pm$ 1.19) 25.63 ($\pm$ 0.00) 27.49 ($\pm$ 0.00)
NeuralODE chen2018neural 3.38 ($\pm$ 0.00) 43.40 ($\pm$ 7.99) -40.37 ($\pm$ 21.82) 53.88 ($\pm$ 0.76) -14.75 ($\pm$ 14.88) 55.36 ($\pm$ 0.90) -36.16 ($\pm$ 12.98) 45.61 ($\pm$ 11.22) -83.55 ($\pm$ 10.01) 32.93 ($\pm$ 18.26) -85.20 ($\pm$ 1.78) 31.35 ($\pm$ 13.08) 38.03 ($\pm$ 14.49)
HigherOrder DMD LeClainche2017 0.94 ($\pm$ 0.00) 51.77 ($\pm$ 0.00) -84.40 ($\pm$ 0.00) 54.88 ($\pm$ 0.00) -90.53 ($\pm$ 0.00) 56.51 ($\pm$ 0.00) -90.80 ($\pm$ 0.00) 66.85 ($\pm$ 0.00) -81.60 ($\pm$ 0.00) 49.74 ($\pm$ 0.00) -11.33 ($\pm$ 0.00) 59.04 ($\pm$ 0.00) 31.22 ($\pm$ 0.00)
Baseline Average -4.73 ($\pm$ 0.00) 51.71 ($\pm$ 0.00) -91.20 ($\pm$ 0.00) 54.88 ($\pm$ 0.00) -91.87 ($\pm$ 0.00) 56.50 ($\pm$ 0.00) -91.33 ($\pm$ 0.00) 65.97 ($\pm$ 0.00) -91.07 ($\pm$ 0.00) 51.93 ($\pm$ 0.00) -90.27 ($\pm$ 0.00) 57.08 ($\pm$ 0.00) 60.88 ($\pm$ 0.00)
Baseline Zeros -39.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -93.33 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -93.47 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -93.73 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -93.73 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -93.73 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00)
(a) Model Scores on Lorenz Dataset
Model Avg Score E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12
Reservoir jaeger_echo_no_date; maass_computational_2004; pathak_model-free_2018 18.88 ($\pm$ 0.00) 99.97 ($\pm$ 0.00) 88.78 ($\pm$ 0.80) 88.61 ($\pm$ 0.04) 23.47 ($\pm$ 4.47) 80.73 ($\pm$ 0.06) -2.57 ($\pm$ 13.48) -12.38 ($\pm$ 6.39) -12.56 ($\pm$ 24.09) -100.00 ($\pm$ 30.40) -100.00 ($\pm$ 100.00) 32.39 ($\pm$ 4.16) 40.08 ($\pm$ 3.93)
LSTM hochreiter1997lstm 15.61 ($\pm$ 0.00) 95.22 ($\pm$ 0.61) -1.88 ($\pm$ 14.14) 90.11 ($\pm$ 0.01) -43.39 ($\pm$ 59.37) 79.83 ($\pm$ 0.07) -27.46 ($\pm$ 35.57) 7.28 ($\pm$ 1.75) 48.74 ($\pm$ 2.21) 4.45 ($\pm$ 7.68) 28.81 ($\pm$ 18.34) -54.07 ($\pm$ 0.00) -40.31 ($\pm$ 0.00)
ODE-LSTM coelho2024odelstm 15.28 ($\pm$ 0.00) 80.09 ($\pm$ 0.27) 0.48 ($\pm$ 50.76) 88.65 ($\pm$ 0.06) -31.46 ($\pm$ 22.33) 52.18 ($\pm$ 0.42) -5.86 ($\pm$ 11.52) 1.71 ($\pm$ 7.42) 49.55 ($\pm$ 7.15) 6.37 ($\pm$ 3.63) 8.52 ($\pm$ 3.08) -54.07 ($\pm$ 0.00) -12.76 ($\pm$ 24.73)
DeepONet deeponet 6.99 ($\pm$ 0.00) 36.52 ($\pm$ 3.85) 17.41 ($\pm$ 6.82) -1.45 ($\pm$ 17.51) 6.52 ($\pm$ 3.27) 6.29 ($\pm$ 1.42) 24.50 ($\pm$ 3.42) -9.48 ($\pm$ 4.27) 1.49 ($\pm$ 1.46) -1.93 ($\pm$ 0.00) -0.15 ($\pm$ 0.00) -4.60 ($\pm$ 7.22) 8.77 ($\pm$ 4.07)
KAN liu2025kankolmogorovarnoldnetworks 6.54 ($\pm$ 0.00) -4.43 ($\pm$ 1.11) 4.89 ($\pm$ 0.80) 50.36 ($\pm$ 0.86) 5.29 ($\pm$ 1.52) 36.93 ($\pm$ 1.20) 24.69 ($\pm$ 8.20) -22.46 ($\pm$ 13.79) 26.47 ($\pm$ 15.12) -43.06 ($\pm$ 15.77) 1.75 ($\pm$ 11.01) 0.83 ($\pm$ 0.17) -2.75 ($\pm$ 2.67)
Baseline Zeros 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00)
Baseline Average -3.02 ($\pm$ 0.00) -3.39 ($\pm$ 0.00) 4.03 ($\pm$ 0.00) 0.01 ($\pm$ 0.00) 0.15 ($\pm$ 0.00) 0.40 ($\pm$ 0.00) 0.17 ($\pm$ 0.00) -9.23 ($\pm$ 0.00) 7.32 ($\pm$ 0.00) -7.12 ($\pm$ 0.00) 13.31 ($\pm$ 0.00) -27.97 ($\pm$ 0.00) -13.88 ($\pm$ 0.00)
SINDy sindy; ensemblesindy -5.23 ($\pm$ 0.00) 22.53 ($\pm$ 0.00) 21.35 ($\pm$ 0.00) -2.91 ($\pm$ 0.00) -100.00 ($\pm$ 0.00) -1.23 ($\pm$ 0.00) -88.53 ($\pm$ 0.00) -0.22 ($\pm$ 0.00) 45.42 ($\pm$ 0.00) -14.00 ($\pm$ 0.00) 34.37 ($\pm$ 0.00) 10.01 ($\pm$ 0.00) 10.51 ($\pm$ 0.00)
OptDMD askham2018variable -11.04 ($\pm$ 0.00) 53.36 ($\pm$ 0.00) -100.00 ($\pm$ 0.04) 6.90 ($\pm$ 0.02) -90.94 ($\pm$ 100.00) 8.82 ($\pm$ 0.17) 19.32 ($\pm$ 4.62) -11.10 ($\pm$ 0.00) 26.41 ($\pm$ 0.00) -71.97 ($\pm$ 28.68) 19.52 ($\pm$ 32.10) 6.00 ($\pm$ 0.00) 1.12 ($\pm$ 0.01)
FNO li2021fourier -18.51 ($\pm$ 0.00) 69.43 ($\pm$ 15.58) -100.00 ($\pm$ 100.00) 17.67 ($\pm$ 35.21) 26.15 ($\pm$ 12.96) 15.85 ($\pm$ 30.12) 16.86 ($\pm$ 23.23) -22.17 ($\pm$ 37.79) -100.00 ($\pm$ 100.00) -46.09 ($\pm$ 17.63) -100.00 ($\pm$ 66.31) 0.76 ($\pm$ 0.00) -0.53 ($\pm$ 0.00)
PyKoopman bruntonkutzkoopmanreview22; Pan2024 -20.11 ($\pm$ 0.00) 14.60 ($\pm$ 0.57) 18.14 ($\pm$ 11.56) 0.00 ($\pm$ 0.00) -100.00 ($\pm$ 55.85) 0.01 ($\pm$ 0.00) -52.27 ($\pm$ 30.45) -17.37 ($\pm$ 3.36) -100.00 ($\pm$ 100.00) -6.90 ($\pm$ 0.00) 0.01 ($\pm$ 0.00) 2.41 ($\pm$ 9.30) 0.02 ($\pm$ 0.06)
HigherOrder DMD LeClainche2017 -24.98 ($\pm$ 0.00) -100.00 ($\pm$ 0.00) -100.00 ($\pm$ 0.00) -0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) -0.18 ($\pm$ 0.28) -100.00 ($\pm$ 100.00) -0.03 ($\pm$ 0.07) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.46 ($\pm$ 0.00)
NeuralODE chen2018neural -33.46 ($\pm$ 0.00) -36.06 ($\pm$ 15.82) 2.99 ($\pm$ 26.81) -100.00 ($\pm$ 11.75) 24.29 ($\pm$ 4.42) -100.00 ($\pm$ 9.21) 20.13 ($\pm$ 25.53) -56.98 ($\pm$ 6.69) -100.00 ($\pm$ 100.00) -98.34 ($\pm$ 8.54) 30.23 ($\pm$ 20.48) 2.05 ($\pm$ 0.23) 10.13 ($\pm$ 0.22)
Spacetime zhang2023spacetime -45.58 ($\pm$ 0.00) 43.49 ($\pm$ 100.00) -100.00 ($\pm$ 100.00) -42.95 ($\pm$ 0.00) 13.70 ($\pm$ 0.00) -43.20 ($\pm$ 0.00) -100.00 ($\pm$ 100.00) -35.56 ($\pm$ 7.78) -100.00 ($\pm$ 100.00) -57.17 ($\pm$ 0.00) -100.00 ($\pm$ 100.00) -31.28 ($\pm$ 0.00) 6.04 ($\pm$ 0.00)
(b) Model Scores on Kuramoto–Sivashinsky Dataset
The training and testing are identical as for the spatio-temporal KS system described above aside from the long range (climate) forecast score. Data matrices for testing and training are of the same form as in Section 2.1 with $n=3$ being the dimension of the dynamical system. Since in this case there is no spatial coordinate it is no longer possible to use the power spectral density of the differential equation to evaluate the long-time performance. Instead, for this system, we evaluate the long-time forecasting based on the distribution of values in the state-space over the last $k$ time steps (e.g. $k=500$). For this we compare the histograms of the distribution of predicted and true solution trajectories in the following way. The histogram for a time series is computed using the histogram command with a set number of bins (e.g., $bins=41$ for our current Lorenz evaluation). The difference of the histogram between the truth ($x,y$ and $z$) and prediction ($\tilde{x},\tilde{y}$ and $\tilde{z}$) for each variable is measured in an $\ell_{1}$-sense:
$s_{{\mbox{\tiny LT}}}(x,\tilde{x})=\frac{\|\text{Hist}_{x}-Hist_{\tilde{x}}\|_{1}}{\|\text{Hist}\|_{1}}.$
From this the long-time error score for the Lorenz system is composed of the distributional error in each coordinate:
$S^{\mathrm{(Lorenz)}}_{{\mbox{\tiny LT}}}({\bf X},\tilde{\bf X})=(s_{{\mbox{\tiny LT}}}(x,\tilde{x})+s_{{\mbox{\tiny LT}}}(y,\tilde{y})+s_{{\mbox{\tiny LT}}}(z,\tilde{z}))/3\hskip 28.90755pt\mbox{(climate forecast)}.$
As with the spatio-temporal system and the power spectral density, this gives a simple measure of the accuracy of the prediction from a statistical viewpoint since long-time prediction is well beyond the Lyapunov time which would not allow for a least-square match between trajectories of the truth and prediction.
2.3 Composite Score
We compute a composite score $\bar{E}$ per dataset from metrics $E_{1}$ through $E_{12}$ by averaging the resulting scores for each method. This score is evaluated per method, not per model. Thus, each method can fit a model for each task and produce the best possible score. All scores are clipped such that $E_{i}\in[-100,100]$, thus $\bar{E}\in[-100,100]$. Methods that cannot produce a result for a given task receive the minimum score 100-100.
3 Methods, Baselines and Results
We characterized twelve highly-cited modeling methods on our ctf4science datasets. Table 1 shows all scored methods and their resulting performance scores. For details on the scored methods, as well as the hyperparameter tuning and evaluation procedures, please refer to the appendix. In addition, we also provide the scores of six zero-shot time-series forecasting foundation models in Table 5 of the appendix. The ctf4science includes two naive baseline methods: predicting zero and predicting the average. In our evaluations, we use average prediction as the baseline for the Lorenz dataset and zero prediction as the reference baseline for KS dataset.
In Fig. 3, we show all evaluated methods per dataset including the naive baselines—constant and average—ranked by their $\bar{E}$. The difference in dimensionality, dynamics, and long-term trajectory stability between Lorenz and KS results in radically different performance distributions. Further, while some models score high on specific tasks, no model scores high-across all tasks (see Table 1). Overall, the results demonstrate that each dataset and task is challenging enough to produce a distribution of scores that characterizes the methods.
Figure 4: Top three performing models per metric on the (a) Lorenz and (b) KS dataset. The blue baseline line here corresponds to the constant zero prediction. This baseline is not producing a score of zero in long-time predictions for the Lorenz dataset due to the different long-time evaluation methods used for KS and Lorenz. KS uses spectral $L_{2}$-error whereas Lorenz uses histogram $L_{2}$-error.
A complete overview of all model’s performance metrics on the Lorenz dataset can be found in table 1(a). The overall score performance for each method in in Fig. 3 while the top three performers in each error category are shown is shown in Fig. 4(a). A complete overview of all model’s performance metrics on the KS dataset can be found in table 1(b). The overall score performance for each method in in Fig. 3 while the top three performers in each error category are shown is shown in Fig. 4(b).
3.1 Observations
Applying the "ImageNet recipe" (fixed public data, objective metrics, leaderboarded methods) to dynamic systems poses new challenges. Scientific models are not trivial to compare, as they range from assumption-rich, high-fidelity approaches to generic, assumption-free, data-hungry models. While the low-dimensional chaotic Lorenz ODE is canonical, easy to synthesize, and analytically transparent, it is chaotic. Chaos guarantees that any forecaster—even the ground-truth solver—accumulates exponential error beyond 3 Lyapunov times, so "predict-the-mean" becomes the rational long-horizon baseline.
Methods therefore succeed or fail depending on whether their implicit assumptions match the task: SINDy excels when its candidate library contains the true terms; operator learners and PINNs might under-perform because they were designed for smooth function-to-function or interpolation problems, not autoregressive time marching; generic RNN-style models struggle at the low data limit, while reservoir models are very well adapted for chaotic time series. Simultaneously, we also see some methods unexpectedly outperformed others in contexts they were not designed for (e.g., DeepONet applied to an autoregressive task on temporal, rather than spatio-temporal data). In essence, ctf4science works as intended. Every task-dataset combination acts as a search light illuminating the performance space within which modeling methods exist and provide insight into which method can tackle which under which conditions.
We begin by presenting a ranking of all methods evaluated from their composite score (See Fig. 3 and Table 1). We present the top 3 models and the constant prediction baseline for each metric from $E1$ through $E12$. The results highlight how the diversity of methods developed have definitive strengths and weaknesses on the various tasks. Thus depending on the task, the appropriate method should be deployed. The CTF provides the critical evaluation metrics necessary for making such decisions.
4 Limitations & Future Work
We are launching ctf4science in a limited scope with three datasets: a dynamical system (Lorenz) and two spatio-temporal system (KS and SST). The evaluation metrics test short- and long-time forecasting and reconstruction under the challenges of noise, limited data and parametric dependency. There are many more datasets and tasks that could and should be considered for science and engineering, most notably tasks in control. This CTF is an important first step to establish fair comparisons among modeling methods on truly withheld test sets. In future versions, more challenging datasets, real world datasets, and more tasks, including control tasks will be integrated.
A key limiting factor in achieving high-scores on the current CTF datasets is the small dataset size, which hamstrings large machine learning models from performing at their best. This was by design, since in many engineering systems, limited data availability is a practical reality. We will expand our collection of datasets and scoring metrics to larger datasets in the future.
Furthermore, the current selection of models is only a starting point. We fully expect that extensions to standard methods could outperform our results (e.g. PINNswang2024respecting). We want to improve on the current results together with the broader research community. ctf4science will help us find successful variations and new applications to existing methods.
While wall-clock time is a useful metric for assessing the potential speed advantage of ML methods over traditional approachesMcGreivy2024, our focus here is on evaluating model suitability for certain tasks. Wall-clock time depends on factors such as hardware configuration, implementation, parallelization, and library efficiency. Nevertheless, we provide our time measurement of each model’s training and evaluation pipeline in the appendix (Table 4) as a rough indication of computational burden.
5 Conclusion
We developed a CTF that scores modeling approaches on a diversity of tasks that are prototypical in science and engineering. The canonical Lorenz and KS systems form an accepted testbench for demonstrating the effectiveness of modeling methods in scientific machine learning literature and act as the starting point of our benchmark. Our work builds a fair and multi-dimensional comparison between methods that is based on a true hidden test set—limiting the risk of "hacked" scores.
CTFs have transformed the research fields that embraced them, such as computer vision, speech and language processing. CTFs have also been critical in identifying protein structure from sequence Kryshtafovych_CASP, leading to the Nobel Prize in Chemistry. Scientific machine learning is now mature enough as a field that a CTF is warranted and needed in order to fairly and accurately evaluate emerging algorithms, especially on the diversity of tasks critical to science and engineering. This work marks the beginning of a sustained effort to provide a neutral and fair comparison between methods and tasks, and thereby boost transparency and competition in machine learning for science.
The central tension our experiment exposes is that scientific ML methods live on a spectrum from assumption-rich, high fidelity to generic, assumption-free, data-hungry models. We see the present CTF as the microscope slide on which this spectrum first becomes visible. Our roadmap adds diverse systems (non-chaotic ODEs, PDEs, stochastic SDEs, experimental datasets), multiple task types (forecasting, system identification, imputation, control), and configuration files that declare what priors each submission may exploit. By exposing where and why celebrated learning algorithms misalign with specific scientific goals, the current CTF is not a verdict on their value but an invitation to researchers in the community to refine architectures and to co-create a truly comprehensive benchmark suite for scientific machine learning; enabling the discovery of scientific breakthroughs and foundational world models.
Acknowledgments and Disclosure of Funding
The authors acknowledge support from the National Science Foundation AI Institute in Dynamic Systems (grant number 2112085). GM acknowledges support from the EPSRC programme grant in ‘The Mathematics of Deep Learning’ (project EP/V026259/1). The Hyperparameter tuning and final evaluation of all models were carried out on the Dutch national supercomputer Snellius, provided by SURF.
Appendix A Appendix
This document contains the supplementary materials for the Common Task Framework For a Critical Evaluation of Scientific Machine Learning Algorithms paper. For each model that was evaluated on the CTF4Science, we share additional implementation and hyperparameter tuning details. This document assumes familiarity with the main text and thus does not redefine terms and details covered in the main text, such as the scoring metrics E1E12E1-E12.
A.1 Dataset Files and Evaluation Metrics
Table 2: Files and corresponding evaluation metrics (E1–E12) for benchmark datasets.
Score Test Task Train / Burn-in File(s) Ground Truth File
E1 Forecasting Short-time $\mathbf{X}_{1\text{train}}$ $\mathbf{X}_{1\text{test}}$
E2 Forecasting Long-time $\mathbf{X}_{1\text{train}}$ $\mathbf{X}_{1\text{test}}$
E3 Noisy (medium) Reconstruction (denoising) $\mathbf{X}_{2\text{train}}$ $\mathbf{X}_{2\text{test}}$
E4 Noisy (medium) Forecast (long-time) $\mathbf{X}_{2\text{train}}$ $\mathbf{X}_{3\text{test}}$
E5 Noisy (high) Reconstruction (denoising) $\mathbf{X}_{3\text{train}}$ $\mathbf{X}_{4\text{test}}$
E6 Noisy (high) Forecast (long-time) $\mathbf{X}_{3\text{train}}$ $\mathbf{X}_{5\text{test}}$
E7 Limited Data (clean) Forecast (short-time) $\mathbf{X}_{4\text{train}}$ $\mathbf{X}_{6\text{test}}$
E8 Limited Data (clean) Forecast (long-time) $\mathbf{X}_{4\text{train}}$ $\mathbf{X}_{6\text{test}}$
E9 Limited Data (noisy) Forecast (short-time) $\mathbf{X}_{5\text{train}}$ $\mathbf{X}_{7\text{test}}$
E10 Limited Data (noisy) Forecast (long-time) $\mathbf{X}_{5\text{train}}$ $\mathbf{X}_{7\text{test}}$
E11 Parametric Generalization Interpolation forecast $\mathbf{X}_{6,7,8\text{train}}$ / $\mathbf{X}_{9\text{train}}$ $\mathbf{X}_{8\text{test}}$
E12 Parametric Generalization Extrapolation forecast $\mathbf{X}_{6,7,8\text{train}}$ / $\mathbf{X}_{10\text{train}}$ $\mathbf{X}_{9\text{test}}$
Table 3: Matrix shapes and indices for the Lorenz dataset (left) and Kuramoto-Sivashinsky dataset (right). Start and end index refer to relative time-steps in the simulation used to generate the dataset matrices. Each successive index represents one $\Delta t$ time-step.
Lorenz
Matrix Shape Start Index End Index
$\mathbf{X}_{1\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{2\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{3\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{4\text{train}}$ $[100,3]$ 0 $100$
$\mathbf{X}_{5\text{train}}$ $[100,3]$ 0 $100$
$\mathbf{X}_{6\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{7\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{8\text{train}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{9\text{train}}$ $[100,3]$ $9900$ $10000$
$\mathbf{X}_{10\text{train}}$ $[100,3]$ $9900$ $10000$
$\mathbf{X}_{1\text{test}}$ $[1000,3]$ $10000$ $11000$
$\mathbf{X}_{2\text{test}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{3\text{test}}$ $[1000,3]$ $10000$ $11000$
$\mathbf{X}_{4\text{test}}$ $[10000,3]$ 0 $10000$
$\mathbf{X}_{5\text{test}}$ $[1000,3]$ $10000$ $11000$
$\mathbf{X}_{6\text{test}}$ $[1000,3]$ $100$ $1100$
$\mathbf{X}_{7\text{test}}$ $[1000,3]$ $100$ $1100$
$\mathbf{X}_{8\text{test}}$ $[1000,3]$ $10000$ $11000$
$\mathbf{X}_{9\text{test}}$ $[1000,3]$ $10000$ $11000$
Kuramoto-Sivashinsky
Matrix Shape Start Index End Index
$\mathbf{X}_{1\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{2\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{3\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{4\text{train}}$ $[100,1024]$ 0 $100$
$\mathbf{X}_{5\text{train}}$ $[100,1024]$ 0 $100$
$\mathbf{X}_{6\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{7\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{8\text{train}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{9\text{train}}$ $[100,1024]$ $9900$ $10000$
$\mathbf{X}_{10\text{train}}$ $[100,1024]$ $9900$ $10000$
$\mathbf{X}_{1\text{test}}$ $[1000,1024]$ $10000$ $11000$
$\mathbf{X}_{2\text{test}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{3\text{test}}$ $[1000,1024]$ $10000$ $11000$
$\mathbf{X}_{4\text{test}}$ $[10000,1024]$ 0 $10000$
$\mathbf{X}_{5\text{test}}$ $[1000,1024]$ $10000$ $11000$
$\mathbf{X}_{6\text{test}}$ $[1000,1024]$ $100$ $1100$
$\mathbf{X}_{7\text{test}}$ $[1000,1024]$ $100$ $1100$
$\mathbf{X}_{8\text{test}}$ $[1000,1024]$ $10000$ $11000$
$\mathbf{X}_{9\text{test}}$ $[1000,1024]$ $10000$ $11000$
Table 4: Average model performances for each metric group on each dataset. E1-E6 demonstrate reconstruction and forecasting performance, E7-E10 demonstrate low-data regime performance, and E11-E12 show parametric generalization performance.
Model E1-E6 E7-E10 E11-E12
Baseline Zeros -46.76 ($\pm$ 0.00) -46.87 ($\pm$ 0.00) 0.00 ($\pm$ 0.00)
Baseline Average -18.55 ($\pm$ 0.00) -15.86 ($\pm$ 0.00) 58.98 ($\pm$ 0.00)
Reservoir [jaeger_echo_no_date, maass_computational_2004, pathak_model-free_2018] 55.77 ($\pm$ 21.25) 31.01 ($\pm$ 8.59) 99.89 ($\pm$ 0.06)
KAN [liu2025kankolmogorovarnoldnetworks] 45.17 ($\pm$ 20.20) 12.57 ($\pm$ 21.60) 51.27 ($\pm$ 2.03)
HigherOrder DMD [LeClainche2017] -17.10 ($\pm$ 0.00) 5.91 ($\pm$ 0.00) 45.13 ($\pm$ 0.00)
OptDMD [askham2018variable] 5.72 ($\pm$ 0.00) 23.55 ($\pm$ 0.00) 59.46 ($\pm$ 0.00)
PyKoopman [bruntonkutzkoopmanreview22, Pan2024] 32.94 ($\pm$ 0.12) -37.70 ($\pm$ 0.54) 26.56 ($\pm$ 0.00)
LSTM [hochreiter1997lstm] 78.07 ($\pm$ 7.44) 41.33 ($\pm$ 12.60) 70.34 ($\pm$ 0.00)
ODE-LSTM [coelho2024odelstm] 49.46 ($\pm$ 8.66) 30.60 ($\pm$ 15.78) 40.42 ($\pm$ 0.00)
Spacetime [zhang2023spacetime] 42.05 ($\pm$ 18.00) 21.27 ($\pm$ 8.66) 65.70 ($\pm$ 0.00)
DeepONet [deeponet] 68.19 ($\pm$ 15.02) 30.57 ($\pm$ 9.34) 81.10 ($\pm$ 9.77)
SINDy [sindy, ensemblesindy] 29.52 ($\pm$ 0.46) 30.60 ($\pm$ 15.78) 48.73 ($\pm$ 0.00)
FNO [li2021fourier] 25.70 ($\pm$ 31.14) 1.18 ($\pm$ 35.58) 43.62 ($\pm$ 9.88)
NeuralODE [chen2018neural] 10.23 ($\pm$ 9.89) -22.55 ($\pm$ 10.32) 34.69 ($\pm$ 13.78)
(a) Average model performances for each metric group on Lorenz Dataset
Model E1-E6 E7-E10 E11-E12
Baseline Zeros 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 0.00 ($\pm$ 0.00)
Baseline Average 0.23 ($\pm$ 0.00) 1.07 ($\pm$ 0.00) -20.92 ($\pm$ 0.00)
Reservoir [jaeger_echo_no_date, maass_computational_2004, pathak_model-free_2018] 63.16 ($\pm$ 3.14) -56.24 ($\pm$ 40.22) 36.23 ($\pm$ 4.04)
KAN [liu2025kankolmogorovarnoldnetworks] 19.62 ($\pm$ 2.28) -9.33 ($\pm$ 13.92) -0.96 ($\pm$ 1.42)
HigherOrder DMD [LeClainche2017] -33.33 ($\pm$ 0.00) -25.05 ($\pm$ 25.09) 0.23 ($\pm$ 0.00)
OptDMD [askham2018variable] -17.09 ($\pm$ 17.48) -9.28 ($\pm$ 15.19) 3.56 ($\pm$ 0.01)
PyKoopman [bruntonkutzkoopmanreview22, Pan2024] -19.92 ($\pm$ 16.40) -31.07 ($\pm$ 25.84) 1.21 ($\pm$ 4.68)
LSTM [hochreiter1997lstm] 32.07 ($\pm$ 18.29) 22.32 ($\pm$ 7.49) -47.19 ($\pm$ 0.00)
ODE-LSTM [coelho2024odelstm] 30.68 ($\pm$ 14.23) 16.54 ($\pm$ 5.32) -33.42 ($\pm$ 12.36)
Spacetime [zhang2023spacetime] -38.16 ($\pm$ 50.00) -73.18 ($\pm$ 51.95) -12.62 ($\pm$ 0.00)
DeepONet [deeponet] 14.96 ($\pm$ 6.05) -2.52 ($\pm$ 1.43) 2.08 ($\pm$ 5.65)
SINDy [sindy, ensemblesindy] -24.80 ($\pm$ 0.00) 16.39 ($\pm$ 0.00) 10.26 ($\pm$ 0.00)
FNO [li2021fourier] 7.66 ($\pm$ 36.18) -67.06 ($\pm$ 55.43) 0.11 ($\pm$ 0.00)
NeuralODE [chen2018neural] -31.44 ($\pm$ 15.59) -56.27 ($\pm$ 33.93) 6.09 ($\pm$ 0.22)
(b) Average model performances for each metric group on KS Dataset
A.2 Evaluations
A.2.1 Hyperparameter Optimization
Hyperparameter optimization is performed in our ctf4science Python package [^3] Available at https://github.com/CTF-for-Science/ctf4science using the tune_module.py script. We employ Ray Tune [liaw2018raytune] for systematic hyperparameter optimization across all models. Hyperparameters are defined in YAML configuration files specifying parameter types, bounds, and sampling distributions. Multiple parameter types are supported, including continuous distributions (uniform, log-uniform), discrete distributions (random integer, log-random integer), and categorical choices.
The optimization follows a trial-based approach where each trial randomly samples a hyperparameter configuration from the defined search space. Each trial trains the model using a train/validation split of the original training dataset. The tune_module.py script splits the training data into train and validation sets, using the latter exclusively for evaluation. Thus, the test set remains unseen during hyperparameter tuning.
Optimization terminates when either a predefined number of trials or a time budget is reached. We employ ASHA (Asynchronous Successive Halving Algorithm) scheduling [li2020asha] for early stopping of poorly performing trials. Resource allocation is automatically managed, distributing trials across available computational resources.
For our results, each combination of model, dataset, and pair_id is allocated 8 hours of tuning time on dedicated nodes equipped with 1 NVIDIA A100 GPU with 40 GiB VRAM and 18 CPU cores from an Intel Xeon Platinum 8360Y processor with 120GiB RAM. Some models complete tuning in less than the alotted time.
A.2.2 Evaluation
Model evaluation is performed using our ctf4science Python package [^4] Available at https://github.com/CTF-for-Science/ctf4science’s benchmark_module.py script. Once hyperparameter tuning is complete, the best-performing parameters on the validation set are used to retrain the model on the full training dataset. The retraining and subsequent evaluation on the test dataset are repeated five times, using different random seeds where possible. We report the mean and standard deviation of the resulting scores across these five runs as indicators of model stability. For models that do not rely on random seeds, the standard deviation is zero. Reported standard deviation values are clipped to a maximum of 100.
A.2.3 Wall-Clock Time
McGreivy and Hakim [McGreivy2024] compared ML methods with traditional approaches under conditions of either equal accuracy or equal runtime, motivated by the claims of the methods in their study that those methods achieve comparable accuracy with improved computational efficiency. In contrast, we take a step back to first examine whether ML methods can achieve reasonable accuracy at all. Therefore, our focus is on the accuracy metrics designed in the paper. Although our goal is not to provide a fair assessment of the speed gain of the ML methods, we nevertheless report the computational costs of the individual models in their current implementations for context. Wall-clock time is measured by our ctf4science package’s performance_module.py script. The total wall-clock time, in seconds, required to train and evaluate each model via our package’s run.py scripts without the visualization option is provided in Table 5.
Table 5: Model mean wall clock times for each pair_id on each dataset
Model pair_id 1 pair_id 2 pair_id 3 pair_id 4 pair_id 5 pair_id 6 pair_id 7 pair_id 8 pair_id 9
Baseline Zeros 0 0 0 0 0 0 0 0 0
Baseline Average 0 0 0 0 0 0 0 0 0
Reservoir [jaeger_echo_no_date, maass_computational_2004, pathak_model-free_2018] 2 12 6 17 2 1 1 17 18
KAN [liu2025kankolmogorovarnoldnetworks] 186 134 180 25 1498 88 85 346 377
HighOrder DMD [LeClainche2017] 0 0 0 0 0 0 0 0 0
OptDMD [askham2018variable] 4 5 5 3 3 0 0 0 0
PyKoopman [bruntonkutzkoopmanreview22, Pan2024] 0 0 0 0 1 0 0 0 0
LSTM [hochreiter1997lstm] 1377 2723 146 2154 1293 51 54 689 485
ODE-LSTM [coelho2024odelstm] 15667 15876 12234 15057 14517 231 172 14447 15073
Spacetime [zhang2023spacetime] 331 832 469 1187 1035 28 27 847 744
DeepONet [deeponet] 234 2 290 39 57 39 40 59 87
SINDy [sindy, ensemblesindy] 1080 937 2745 3 72 189 70 153 248
FNO [li2021fourier] 417 1098 924 1477 375 19 21 907 2184
NeuralODE [chen2018neural] 9468 2172 848 2390 786 51 27 4460 3589
PINN [raissi2019physics] 77 77 76 76 76 76 76 76 76
(a) Mean Wall Clock Times on Lorenz Dataset in Seconds
Model pair_id 1 pair_id 2 pair_id 3 pair_id 4 pair_id 5 pair_id 6 pair_id 7 pair_id 8 pair_id 9
Baseline Zeros 0 0 0 0 0 0 0 0 0
Baseline Average 0 0 0 0 0 0 0 0 0
Reservoir [jaeger_echo_no_date, maass_computational_2004, pathak_model-free_2018] 306 424 637 185 107 28 26 64 245
KAN [liu2025kankolmogorovarnoldnetworks] 1367 77 1797 159 1495 2406 1851 2286 1840
HigherOrder DMD [LeClainche2017] 2 4 2 3 2 0 1 4 5
OptDMD [askham2018variable] 78 77 89 57 46 1 1 11 15
PyKoopman [bruntonkutzkoopmanreview22, Pan2024] 44 2 45 3 62 1 0 16 3
LSTM [hochreiter1997lstm] 3243 369 $14$ 835 728 50 50 1830 1171
ODE-LSTM [coelho2024odelstm] 22067 2270 2506 21957 17956 375 282 17238 1535
Spacetime [zhang2023spacetime] 6611 13981 1952 9439 6715 19 22 1110 3280
DeepONet [deeponet] 1348 118 2414 334 2817 160 36 1965 6272
SINDy [sindy, ensemblesindy] 53950 157 9 24 6731 139 649 16 348
FNO [li2021fourier] 762 930 2154 597 2877 17 10 2852 30
NeuralODE [chen2018neural] 2841 1635 421 451 196 39 21.24 4528 2957.52
(b) Mean Wall Clock Times on Kuramoto–Sivashinsky Dataset in Seconds
A.3 Foundation Model Results
We evaluated the performance of several widely used foundation models on our CTF. Each of these models is advertised as being capable of performing zero-shot time-series forecasting. The results are presented in Table 6. As the foundation models are pre-trained, we did not perform hyperparameter tuning or training. Instead, we provide their one-shot results, reflecting how such models would typically be used in real-world applications.
Table 6: Foundation model performances for each metric on each dataset
Model avg_score E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12
Panda [lai2025panda] -59.60 -69.13 -38.51 -100.00 -38.21 -100.00 -41.20 -97.19 -36.21 -51.01 -35.09 -56.99 -51.60
Moirai [liu2024moiraimoeempoweringtimeseries] -12.07 49.96 -88.53 29.74 -84.33 25.61 -84.67 55.25 -87.20 52.28 -90.73 50.06 27.75
Chronos [ansari2024chronoslearninglanguagetime] -7.27 34.80 -84.67 52.85 -86.53 53.40 -88.00 44.18 -88.47 54.01 -85.13 49.24 57.04
TabPFN [hoo2025tablestimetabpfnv2outperforms] 28.80 51.35 -26.27 84.06 -26.80 79.02 -14.27 31.49 58.00 28.85 27.60 22.54 29.96
LLMTime [gruver2023large] -36.89 4.59 -91.40 0.59 -100.00 0.44 -94.47 4.34 -93.73 4.10 -94.47 8.38 8.99
Sundial [liu2025sundialfamilyhighlycapable] 45.26 53.24 40.30 50.94 39.68 45.32 34.94 45.19 42.04 52.19 44.95 47.37 47.01
(a) Model Scores on Lorenz Dataset
Model avg_score E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12
Panda [lai2025panda] -96.14 -6.28 -100.00 -100.00 -100.00 -100.00 -100.00 -171.75 -100.00 -100.00 -100.00 -103.81 -71.78
Moirai [liu2024moiraimoeempoweringtimeseries] -93.79 -100.00 -100.00 -25.53 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00
Chronos [ansari2024chronoslearninglanguagetime] -23.03 37.89 26.91 -100.00 -6.24 -100.00 3.44 -4.11 0.21 -23.40 -100.00 -7.02 -4.08
TabPFN [hoo2025tablestimetabpfnv2outperforms] -2.51 97.91 3.65 -100.00 2.01 -100.00 1.17 3.66 30.91 -32.50 24.74 12.67 25.70
LLMTime [gruver2023large] -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00 -100.00
Sundial [liu2025sundialfamilyhighlycapable] -0.64 7.17 8.22 4.19 -6.42 -0.75 -3.34 -1.37 1.07 0.74 0.52 -16.28 -1.40
(b) Model Scores on Kuramoto–Sivashinsky Dataset
A.4 Models
A.4.1 Baselines
We implement two baseline models. One of the baselines predicts all zeros. The other baseline predicts the average of the input data per spatial dimension. We do not perform hyperparameter optimization for either of these models.
A.4.2 LSTM/ODE-LSTM
LSTM networks are a specialized type of recurrent neural network (RNN) designed to address the vanishing gradient problem inherent in traditional RNNs [hochreiter1997lstm]. They achieve this through a unique architecture featuring memory cells and gating mechanisms (input, forget, and output gates), which regulate the flow of information over time. These gates enable LSTMs to selectively retain or discard historical data, making them particularly adept at capturing long-term dependencies in sequential data. In time-series forecasting, LSTMs excel at modeling temporal patterns, such as trends, seasonality, and irregular fluctuations, by leveraging past observations to predict future values. Their ability to handle complex, non-linear relationships and variable-length input sequences makes them a robust choice for tasks like stock prediction, energy load forecasting, or weather modeling, where historical context is critical to accurate predictions.
ODE-LSTMs are a flavor of LSTMs that try to further tackle the vanishing gradient problem by using an ODE solver to model the hidden state of the LSTM [coelho2024odelstm]. They show that traditional LSTMs can still suffer from a vanishing or exploding gradient and provide theory demonstrating ODE-LSTMs do not suffer from either of these problems.
We evaluate both a classical LSTM and the ODE-LSTM by searching over the following hyperparameters: hidden_state_size (dimension of the latent space), seq_length (input sequence length), and lr (learning rate).
hyperparameter type min (or options) max (or none)
hidden_state_size randint 3 32
seq_length randint 5 512
lr log_uniform $10^{-5}$ $10^{-2}$
Table 7: Hyperparameter search space for the ODE-LSTM and LSTM models on metrics $E_{1}$ through $E_{6}$ for Lorenz. We train with a batch size of 128 for 200 epochs.
hyperparameter type min (or options) max (or none)
hidden_state_size randint 8 256
seq_length randint 5 512
lr log_uniform $10^{-5}$ $10^{-2}$
Table 8: Hyperparameter search space for the ODE-LSTM and LSTM models on metrics $E_{1}$ through $E_{6}$ for Kuramoto-Sivashinsky. We train with a batch size of 128 for 200 epochs.
hyperparameter type min (or options) max (or none)
hidden_state_size randint 3 32
seq_length randint 5 74
lr log_uniform $10^{-5}$ $10^{-2}$
Table 9: Hyperparameter search space for the ODE-LSTM and LSTM models on metrics $E_{7}$ through $E_{12}$ for Lorenz. We train with a batch size of 5 for $E_{7}$ through $E_{10}$ and a batch size of 128 for $E_{11}$ and $E_{12}$ for 200 epochs.
hyperparameter type min (or options) max (or none)
hidden_state_size randint 8 256
seq_length randint 5 74
lr log_uniform $10^{-5}$ $10^{-2}$
Table 10: Hyperparameter search space for the ODE-LSTM and LSTM models on metrics $E_{7}$ through $E_{12}$ for Kuramoto-Sivashinsky. We train with a batch size of 5 for $E_{7}$ through $E_{10}$ and a batch size of 128 for $E_{11}$ and $E_{12}$ for 200 epochs.
A.4.3 SpaceTime
State-Space Models (SSMs) are mathematical frameworks that describe systems using latent (hidden) states evolving over time, observed through measurable outputs. They are widely used in control theory, signal processing, and time-series analysis to model dynamic systems. Modern adaptations like S4 (Structured State Space for Sequence Modeling) and SpaceTime are deep learning variants of SSMs tailored for sequential data. These models parameterize state transitions with structured matrices to efficiently capture long-range dependencies while remaining computationally tractable. Unlike LSTMs, SSMs are particularly effective at time-series forecasting of long-range dependencies with minimal memory overhead.
SpaceTime [zhang2023spacetime] is one such SSM that claims to be a state-of-the-art model on time-series forecasting and classification tasks. The authors claim that their model captures “complex, long-range, and autoregressive” dependencies, can forecast over long horizons, and is efficient during training and inference. They demonstrate improved performance over the popular S4 SSM and NLinear.
Based on the hyperparameter optimization described in the original paper and the hyperparameters which can be adjusted in the publicly available code, we do a hyperparameter search over the following values: lag (input sequence length), horizon (output sequence length), n_blocks (number of SpaceTime layers in the model encoder), dropout, weight_decay, kernel_dim (dimension of SSM kernel in each block), and lr (learning rate).
hyperparameter type min (or options) max (or none)
lag randint 32 512
horizon randint 32 512
n_blocks choice {3,4,5,6} $\cdot$
dropout choice {0, 0.25} $\cdot$
weight_decay choice {0, 0.0001} $\cdot$
kernel_dim choice {32,64,128} $\cdot$
lr log_uniform $10^{-5}$ $10^{-2}$
Table 11: Hyperparameter search space for the SpaceTime model on metrics $E_{1}$ through $E_{6}$ for Lorenz and Kuramoto-Sivashinsky. We train with a batch size of 128 for 200 epochs.
hyperparameter type min (or options) max (or none)
lag randint 10 45
horizon randint 10 45
n_blocks choice {3,4,5,6} $\cdot$
dropout choice {0, 0.25} $\cdot$
weight_decay choice {0, 0.0001} $\cdot$
kernel_dim choice {32,64,128} $\cdot$
lr log_uniform $10^{-5}$ $10^{-2}$
Table 12: Hyperparameter search space for the SpaceTime model on metrics $E_{7}$ through $E_{10}$ for Lorenz and Kuramoto-Sivashinsky. We train with a batch size of 5 for 200 epochs.
hyperparameter type min (or options) max (or none)
lag randint 10 45
horizon randint 10 45
n_blocks choice {3,4,5,6} $\cdot$
dropout choice {0, 0.25} $\cdot$
weight_decay choice {0, 0.0001} $\cdot$
kernel_dim choice {32,64,128} $\cdot$
lr log_uniform $10^{-5}$ $10^{-2}$
Table 13: Hyperparameter search space for the SpaceTime model on metrics $E_{11}$ through $E_{12}$ for Lorenz and Kuramoto-Sivashinsky. We train with a batch size of 128 for 200 epochs.
A.4.4 Deep Operator Networks
Deep Operator Networks (DeepONets) [deeponet] recently emerged as a powerful tool designed to efficiently model high-dimensional physical systems and complex input-output relationships, as well as to solve challenging problems in scientific machine learning and engineering, such as partial differential equations. Specifically, DeepONets are a class of neural operators which decompose an operator $G:\mathcal{V}\to\mathcal{U}$ between infinite-dimensional functional spaces $\mathcal{V}$ and $\mathcal{U}$ into two cooperating sub-networks, namely branch and trunk net. The trunk encodes the input function $v\in\mathcal{V}:\Omega^{\prime}\subset\mathbb{R}^{d}\to\mathbb{R}^{n_{v}}$ – which is typically sampled at a finite set of $n$ fixed sensors, resulting in the measurement vector $\mathbf{v}\in\mathbb{R}^{n\cdot n_{v}}$ – into $p$ coefficients $\mathbf{b}\left(v\right)\in\mathbb{R}^{p}$. Instead, the branch net provides the evaluation of a neural learnable $p$-dimensional basis $\mathbf{t}\left(\xi\right)\in\mathbb{R}^{p}$ at the spatial coordinates $\xi$ in the domain $\Omega\subset\mathbb{R}^{d}$. Doing so, the value of the output function $u\in\mathcal{U}:\Omega\to\mathbb{R}^{n_{u}}$ at the evaluation point $\xi\in\Omega$ is approximated through the basis expansion
$u(\xi)=G\left(v\right)\left(\xi\right)\approx\mathbf{b}\left(v\right)\cdot\mathbf{t}\left(\xi\right).$
See [deeponet, deeponetapproxtheory, pod-deeponet] for a complete presentation of DeepONets, including also universal approximation theorems for operators. A graphical summary of the DeepONet architecture is available in Figure 5.
Figure 5: Architecture of the Deep Operator Network. The target field at the evaluation point $\xi$ is approximated by the inner product of the outputs of the branch net, which takes as input the measurements $\mathbf{v}$ of the input function $v\in\mathcal{V}$ and returns a set of coefficients $\mathbf{b}(\mathbf{v})$, and the trunk net, which encodes the coordinates $\xi$ into a vector $\mathbf{t}(\xi)$.
DeepONets for dynamical systems
DeepONets are versatile neural architectures designed to learn mappings between functional spaces. DeepONets are traditionally exploited for inferring the space-time evolution of physical variables, such as the solution of partial differential equations, starting from known quantities, such as forcing terms, initial conditions, parameters or control variables [deeponet, pod-deeponet, pi-deeponet-control, latent-deeponet]. However, it is possible to adapt and employ DeepONets in the proposed CTF in order to model and forecast time-series data and dynamical systems, as proposed by, e.g., [forecasting-deeponet, forecasting-deeponet1, forecasting-deeponet2, s-deeponet, s-deeponet1, don-lstm]. Specifically, we consider the operator
$u_{t}(\xi)=G\left(u_{t-1},...,u_{t-k}\right)\left(\xi\right)\approx\mathbf{b}\left(\mathbf{u}_{t-1},...,\mathbf{u}_{t-k}\right)\cdot\mathbf{t}\left(\xi\right)$
where $u_{t}:\Omega\to\mathbb{R}^{n_{u}}$ and $\mathbf{u}_{t}\in\mathbb{R}^{n}$ are, respectively, the solution of the dynamical system under investigation at time $t$ and the corresponding spatial discretization, $k$ is the lag parameter and $\xi\in\Omega\subset\mathbb{R}^{d}$ are the spatial coordinates where to predict the evolution of the dynamics. Along with the evaluation point $\xi$, the trunk input may be enlarged with the time instance $t$ or the time-step $\Delta t$, as proposed by [pod-deeponet, forecasting-deeponet2].
DeepONets implementation
The implementation of DeepONets within the proposed CTF is based on the DeepXDE library [deepxde]. In particular, when dealing with forecasting tasks, we predict the state evolution in an autoregressive manner, and we enlarge the trunk input with the time-step $\Delta t$, as it results in better performance. As proposed by [pod-deeponet], we consider a scaler to normalize the data before training. Moreover, we employ branch and trunk networks with the same number of neurons per hidden layer, so as to reduce the number of hyperparameters.
The Kuramoto-Sivashinsky dataset deals with one-dimensional scalar-valued functions, that is $d=n_{u}=1$. The KS solution is discretized and evaluated at $n=1024$ spatial points uniformly spaced in the domain $\Omega=[0,32\pi]$. Notice that we take into account the same locations across all the input-output pairs, resulting in a lower computational cost.
The Lorenz test case, instead, considers a three-dimensional state variable evolving over time, without spatial dependence. Among different alternatives, we adapt DeepONet in this context by considering the fictitious domain $\Omega={1,2,3}$ and the state function $u_{t}:\Omega=\{1,2,3\}\to\mathbb{R}$ mapping the index $\xi\in\Omega=\{1,2,3\}$ into the $\xi$-th component of the state vector at time $t$. For instance, if $\xi=1$, DeepONet predicts the evolution of the first component of the state variable starting from the past state values encoded by the branch net.
Hyperparameters
The DeepONet hyperparameters mainly concern the neural network architectures and the corresponding training procedure. In addition, the lag parameter determines the length of the past state history fed into the branch input for forecasting. Notice that the lag value cannot be larger than the dimension of burn-in data, and it is set equal to zero when dealing with reconstruction tasks. Table 14 provides a summary of the hyperparameters in play, along with the corresponding search spaces explored for hyperparameters tuning.
hyperparameter type min (or options) max (or none)
lag integer 11 $99$
branch_layers integer 1 5
trunk_layers integer 1 5
neurons integer 1 512
activation choice {"tanh", "relu", "elu"} $\cdot$
initialization choice {"Glorot normal", "He normal"} $\cdot$
optimizer choice { "adam", "L-BFGS" } $\cdot$
learning_rate loguniform $10^{-5}$ $10^{-1}$
epochs integer 10000 10000
Table 14: Hyperparameter search space for DeepONet.
A.4.5 Sparse Identification of Nonlinear Dynamics
Sparse Identification of Nonlinear Dynamics (SINDy) [sindy] is a powerful algorithm designed to discover interpretable and parsimonious governing equations from time-series data. Given the data matrices
$X=\begin{bmatrix}x_{1}(t_{1})&x_{1}(t_{2})&...&x_{1}(t_{m})\\ \vdots&\vdots&\ddots&\vdots\\ x_{n}(t_{1})&x_{n}(t_{2})&...&x_{n}(t_{m})\end{bmatrix};\quad\dot{X}=\begin{bmatrix}\dot{x}_{1}(t_{1})&\dot{x}_{1}(t_{2})&...&\dot{x}_{1}(t_{m})\\ \vdots&\vdots&\ddots&\vdots\\ \dot{x}_{n}(t_{1})&\dot{x}_{n}(t_{2})&...&\dot{x}_{n}(t_{m})\end{bmatrix}$
collecting, respectively, the state vector $\mathbf{x}(t)=[x_{1}(t),...,x_{n}(t)]$ and the corresponding time derivatives $\dot{\mathbf{x}}(t)=[\dot{x}_{1}(t),...,\dot{x}_{n}(t)]$ at the time instances $t_{1},...,t_{m}$, we aim at identifying the (possibly nonlinear) underlying governing equation $\dot{\mathbf{x}}(t)=f(\mathbf{x}(t))$. To this aim, SINDy considers the following approximation
$\dot{X}=\Theta(X)\Xi$
where $\Theta(X)$ is a library of candidate regression terms, such as polynomials or trigonometric functions, while $\Xi$ are the corresponding regression coefficients. Sparsity promoting strategies are crucial to identify simple and interpretable dynamical systems, capable of avoiding overfitting and accurately extrapolating beyond training data. In particular, the regression coefficients $\Xi$ are determined through sparse regression strategies, such as Least Absolute Shrinkage and Selection Operator (LASSO) or Sequentially Thresholded Least SQuares (STLSQ). See Figure 6 for a scheme of the SINDy algorithm on the Lorenz system.
Figure 6: Schematic of the Sparse Identification of Nonlinear Dynamics (SINDy) algorithm from [sindy], demonstrated on the Lorenz equations. The temporal evolution of the state variable and its derivative are collected in the data matrices $X$ and $\dot{X}$. The dynamical system $\dot{X}=\Theta(X)\Xi$ is then identified through sparsity promoting algorithms.
SINDy can easily handle parametric dependencies: indeed, augmenting the state vector with the (possibly time-dependent) parameter values $\boldsymbol{\mu}$ and adding $\boldsymbol{\mu}$-dependent terms in the library $\Theta(X,\boldsymbol{\mu})$, it is possible to identify parametric sparse dynamical systems.
Identifying sparse dynamical systems from high-dimensional data may be computationally expensive. A possible workaround is given by dimensionality reduction techniques, such as Proper Orthogonal Decomposition (POD) [sindy] or autoencoders [sindy-ae], which project state snapshots onto a low-dimensional manifold. SINDy can thus be applied on the low-dimensional latent variables, allowing for efficient and accurate forecasting of the high-dimensional state evolution.
SINDy implementation
The implementation of SINDy is based on the PySINDy library [pysindy]. After collecting the data and approximating the time derivatives through numerical schemes, the SINDy algorithm is applied to identify a sparse dynamical system describing the data evolution over time. The integrator solve_ivp by scipy [scipy] is considered to simulate the system and to predict future state values. Notice that, whenever the identified model is very complex and the integrator fails, the static dynamical system $\dot{\mathbf{x}}=0$ is employed.
The Kuramoto-Sivashinsky dataset deals with the temporal evolution of a chaotic partial differential equation on the spatial domain $0,32\pi]$. The KS solution is discretized and evaluated at $n=1024$ locations, resulting in a collection of high-dimensional snapshots over time. Proper Orthogonal Decomposition (POD) is thus exploited to compress the temporal data, and SINDy is applied to identify the dynamics of the most energetic POD coefficients. Therefore, the KS predictions are retrieved by integrating the SINDy model and projecting the POD coefficients onto the original high-dimensional state space.
Parametric SINDy models are considered when testing the ability of the model to generalize to different parameter values. Since the parameter values employed for data generation are not publicly available, we take into account fictitious values mimicking the interpolatory and extrapolatory regimes.
Hyperparameters
The SINDy algorithm can exploit different differentiation methods to approximate time derivatives, different terms in the library $\Theta(X)$ – such as, e.g., polynomials and/or trigonometric functions up to a chosen order – as well as different sparse regression techniques. Table 15 provides a summary of the hyperparameters in play, along with the corresponding search spaces explored for hyperparameter tuning.
hyperparameter type min (or options) max (or none)
POD_modes integer 11 $50$
differentiation_method choice { "finite_difference", "spline", $\cdot$
"savitzky_golay", "spectral",
"trend_filtered", "kalman" }
differentiation_method_order integer 1 10
feature_library choice { "polynomial", $\cdot$
"Fourier", "mixed" }
feature_library_order integer 11 $10$
optimizer choice {"STLSQ", "SR3", $\cdot$
"SSR", "FROLS"}
threshold choice { "adam", "L-BFGS" } $\cdot$
learning_rate loguniform $10^{-3}$ $10^{3}$
alpha loguniform $10^{-3}$ $10^{1}$
Table 15: Hyperparameter search space for SINDy. The POD_modes parameter has an effect only for the Kuramoto-Sivashinsky test case.
A.4.6 Dynamic Mode Decomposition
The Dynamic Mode Decomposition (DMD) is a data-driven method developed by Schmid [schmid_dynamic_2010] in the fluid dynamics community to identify spatio-temporal coherent structures from high-dimensional data. The DMD algorithm is based on the Singular Value Decomposition (SVD) of a data matrix; in particular, DMD is able to provide a modal decomposition where each mode consists of spatially correlated structures that have the same linear behaviour in time. The DMD method is found to have a significant connection with the Koopman operator theory [rowley_spectral_2009]: in particular, the DMD can be formulated as an algorithm able to learn the best-fit linear dynamical system to advance in time (Figure 7).
There are many variants of DMD, connected to existing techniques from system identification and modal extraction [brunton_data-driven_2022]. Here, we will provide a brief overview of the underlying idea of the original DMD algorithm, from which all the other variants can be derived. The first step is to collect a set of snapshots of the system at different time steps. The data matrix is then constructed by stacking the snapshots in columns, i.e., $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N_{t}}]\in\mathbb{C}^{\mathcal{N}_{h}\times N_{t}}$, where $\mathbf{x}_{k}\in\mathbb{C}^{\mathcal{N}_{h}}$ is the $k$-th snapshot at time $t_{k}$ and $N_{t}$ is the number of snapshots. The original formulation from [schmid_dynamic_2010, rowley_spectral_2009] supposed uniform sampling in time, i.e. $t_{k}=k\Delta t$, where $\Delta t$ is the time step and $t_{k+1}=t_{k}+\Delta t$. Overall, the DMD algorithm seeks the leading spectral decomposition of the best-fit linear operator $\mathbb{A}\in\mathbb{C}^{\mathcal{N}_{h}\times\mathcal{N}_{h}}$ that advances the system in time, i.e.
$\mathbf{x}_{k+1}\approx\mathbb{A}\mathbf{x}_{k}\qquad\longleftrightarrow\qquad\mathbf{X}_{[2:N_{t}]}\approx\mathbb{A}\mathbf{X}_{[1:N_{t}-1]}$
As we said above, the DMD algorithm is based on the SVD of the data matrix $\mathbf{X}$ of rank $r$, which can be written as $\mathbb{X}\simeq\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{*}$: $\mathbf{U}\in\mathbb{C}^{\mathcal{N}_{h}\times r}$ represents the left singular vectors and are also known as modes, describing the dominant spatial structures extracted from the data; the diagonal matrix $\mathbf{\Sigma}\in\mathbb{R}^{r\times r}$ contains the singular values, which are related to the energy/information retained by the modes; in the end, $\mathbf{V}^{*}\in\mathbb{C}^{r\times N_{t}}$ represents the right singular vectors, which are related to the temporal dynamics of the modes. This compression operation allows to compute the dynamical matrix $\mathbb{A}$ in a more efficient way [kutz-dmd_book2016, brunton_data-driven_2022], avoiding the direct inversion of the high-dimensional snapshot matrix.
Figure 7: Scheme of the Dynamic Mode Decomposition algorithm from [kutz-dmd_book2016]. The data matrix $\mathbf{X}$ is constructed by stacking the snapshots in columns. The SVD of the data matrix is computed, and the dynamical matrix is fitted to the data. This allows us to compute the state of the system for future time instances.
Indeed, in the literature different variants of DMD have been proposed: in this context, the High-Order DMD (HODMD) [LeClainche_HoDMD], which exploits time delay embedding to fit the optimal Koopman Operator, and the Optimised DMD (OptDMD) [optdmd2018, bopdmd_2022], which is a variant of DMD that can also use the Bagging algorithm to improve the robustness of the DMD algorithm against noise. This latter variant has been shown to be the most robust and stable algorithm for real-world applications [faraji_data-driven_2024-2]. The implementation of the DMD algorithm is available in the pyDMD package [demo_pydmd_2018, ichinaga_pydmd_2024], which is a Python library for DMD and its variants. The library is designed to be easy to use and flexible, allowing users to customise the algorithm for their specific needs.
Parametric DMD
The extension of DMD to parametric systems is a recent development in the field of system identification. Different approaches have been proposed in the literature; in this work, the implementation of Andreuzzi et al. [andreuzzi_dynamic_2023] within pyDMD is adopted. Up to now, the package does not support the OptDMD algorithm directly, we have implemented a wrapper to use the OptDMD algorithm with the parametric DMD following the same approach of the package, based on the interpolation of the forecasted reduced dynamics. We appreciate that further work and rigorous testing of this implementation are planned for future work. Similar to SINDy, since the parameter values employed for data generation are not publicly available, fictitious values mimicking the interpolatory and extrapolatory regimes have been used.
Hyperparameter tuning
The hyperparameters of the DMD algorithm depend on the specific variant adopted. Every DMD algorithm has a set of hyperparameters that can be tuned to improve the performance of the algorithm; however, the rank of the SVD is common to all of them and plays a crucial role in the reduction process. The HODMD algorithm also includes the delay embedding, defining the size of the lagging window to use. The OptDMD algorithm can also put constraints on the DMD eigenvalues to ensure that the dynamics follow a certain behaviour. In the end, the parametric DMD can operate in two different modes: partitioned and monolithic. The hyperparameters of both DMD algorithms are listed in Tables 16 and 17.
hyperparameter type min (or options) max (or none)
rank randint 33 $50$
delay randint 0 $200$
parametric choice {"partitioned", "monolithic"}
Table 16: Hyperparameter search space for the HODMD algorithm for Lorenz and Kuramoto-Sivashinsky (the parametric hyperparameter has an effect only for metrics $E_{11}$ and $E_{12}$).
hyperparameter type min (or options) max (or none)
rank randint 33 $50$
delay randint 0 $100$
parametric choice { "partitioned", "monolithic"}
eig_constraints choice { "none", "stable", "conjugate_pairs"}
Table 17: Hyperparameter search space for the OptDMD algorithm for Lorenz and Kuramoto-Sivashinsky (the parametric hyperparameter has an effect only for metrics $E_{11}$ and $E_{12}$).
A.4.7 Koopman operator-based dynamic system prediction
The Koopman operator
Koopman operator theory is a useful tool that has found increasing attention in the data-driven scientific computing community and can essentially be seen as an extension of dynamic mode decomposition - viewing the statespace of the dynamic system through the lens of nonlinear observables. This point-of-view dates back to early work by [koopman1931hamiltonian, koopman1932dynamical] and a modern review can be found in [bruntonkutzkoopmanreview22]. We outline the method briefly before describing the set-up for the chosen implementation and our testing on the CTF. Consider a dynamical system (either an ODE or a semi-discretisiation of a PDE) of the form:
$\frac{d\mathbf{x}}{dt}=\mathbf{f}(\mathbf{x}),$
where $\mathbf{f}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ may be a nonlinear forcing. The central idea in Koopman operator theory is then to learn a coordinate transform (i.e. a set of nonlinear observables) $\Phi:\mathbb{R}^{N}\rightarrow\mathbb{R}^{M}$, under which the dynamics becomes (approximately) linear, i.e.
$\frac{d\mathbf{z}}{dt}\approx\mathbf{A}\mathbf{z},\qquad\mathbf{z}(t)=\Phi(\mathbf{x}(t)).$
In this new coordinate system, the exact solution of the linear dynamics is straightforward. The inference of $\mathbf{\Phi}$ and $\mathbf{A}$ can be formulated as a regression problem.
Numerical implementation and parameter choices
In our current CTF test we use the PyKoopman Python library as the main reference point for the Koopman method for dynamic system prediction [Pan2024]. The Python package serves as a good reference since it is regularly maintained and has an up-to-date implementation of several central features of the Koopman operator framework. As mentioned above there are two central parameters that affect the performance of the Koopman method: the observables and the regression method. Exploiting the existing implementation in PyKoopman we allowed in our CTF testing the variation of the following set of parameters:
  • Type of observable: Options include the identity, polynomials of variable degree, time delay (of variable depth), radial basis functions (of variable number) and random Fourier features, as well as the concatation of all of the aforementioned observables with the identity;
  • Type of regressor: DMD, EDMD, HAVOK and KDMD;
  • Regressor rank;
  • Least-squares regularisation and rank of the regularisation (this option is implemented only in EDMD and KDMD).
Note that in principle a neural network-based DMD is also implemented in the PyKoopman package, but in our fine-tuning we found that this lead consistently to worse performance than the above four types of regressors thus we did exclude it from the hyperparameter tuning.
Parametric PyKoopman
Out-of-the-box PyKoopman does not have a parametric implementation, thus in order to test the Koopman method on task 4, we loosely follow [andreuzzi2023dynamic, guo2025learning] and implement a custom parametric version of PyKoopman by spline interpolation of the learned Koopman operator and corresponding eigenfunctions. We acknowledge that further work and rigorous testing of various parametric versions of the Koopman method are required to identify the best performing implementation for task 4.
Further comments on the use with chaotic systems
We note that the performance of the Koopman operator on the KS and Lorenz system is notably subpar, especially when compared to results reported in prior work [pykoopman_documentation]. This is not unexpected and a likely source of challenge is the chaotic nature of both equations, which has also been noticed by the authors of the PyKoopman package. Essentially, in chaotic systems there may not be a dominating low-rank structure that can be learned and exploited with the Koopman method (cf. the section on “Unsuccessful examples of using Dynamic mode decomposition on PDE system” in [pykoopman_documentation]).
Hyperparameter tuning
Based on the available choices implemented in the PyKoopman package and the examples described in the documentation [pykoopman_documentation], we performed a hyperparameter search over the following parameters: type of observable and potential concatenation with the identity, observables integer parameter (representing the polynomial degree in case of polynomial observables, the number of time delay steps in the case of time delay observables and the parameter $D$ in the random Fourier feature case), the number of centers for the radial basis function observables, observables float parameter (representing the radial basis function kernel width and the parameter $\gamma$ in the radial basis function case respectively), regressor type, regressor rank, TLSQ rank (the regularisation rank called only when the regressor is EDMD and KDMD). The details of the parameter space explored are shown in Table 18.
hyperparameter type min (or options) max (or none)
observables choice {Identity, Polynomial, TimeDelay, $\cdot$
RadialBasisFunctions,
RandomFourierFeatures}
Identity concatenation choice {true, false}
Integer parameter randint 1 10
# RBF centers randint 10 1000
Float parameter uniform 0.5 2.0
regressor type choice {DMD,EDMD, HAVOK, KDMD} $\cdot$
regressor rank randint 1 200
TLSQ rank randint 1 200
Table 18: Hyperparameter search space for the PyKoopman model.
A.4.8 Reservoir Computing
In its broadest sense, reservoir computing (RC) is a general machine learning framework for processing sequential data. RC functions by projecting data into a high-dimensional dynamical system and training a simple readout from these dynamics back to a quantity or signal of interest. Although there exists a large and ever-growing body of literature on leveraging physical systems to act as high-dimensional “reservoirs” [tanaka_recent_2019], the most common form of RC remains an echo state network (ESN) [jaeger_echo_2001, maass_real-time_2002]. ESNs are a form of recurrent neural network (RNN) that have been demonstrated to achieve state-of-the-art performance in the forecasting of chaotic dynamical systems [platt_systematic_2022, vlachas_backpropagation_2020]. We now introduce the specific form of ESN we use in evaluating performance on the CTF datasets, following many of the conventions presented in [platt_systematic_2022].
ESNs for Lorenz63 system.
Given a time series $u_{0},\dots,u_{T}$, a randomly instantiated, high-dimensional dynamical system is evolved according to
$h_{t+1}=(1-\alpha)h_{t}+\alpha\tanh\left(W_{hh}h_{t}+W_{hu}u_{t}+\sigma_{b}\mathbf{1}\right)$ (5)
where $\alpha$ is the so-called leak rate hyperparameter, $W_{hh}$ and $W_{hu}$ are fixed, random matrices, $\sigma_{b}$ is a bias hyperparameter and $\mathbf{1}$ denotes a vector of ones. $W_{hh}\in\mathbb{R}^{N_{h}\times N_{h}}$ ($N_{h}$ denotes the number of entries in $h$) is taken to be a random, sparse matrix with density $\approx 2\%$ and non-zero entries sampled from $\mathcal{U}(-1,1)$ and then scaled such that the spectral radius of $W_{hh}$ is $\rho.$ $W_{hu}\in\mathbb{R}^{N_{h}\times N_{u}}$ ($N_{u}$ denotes the number of entries in $u$) is a random matrix with each entry drawn independently from $\mathcal{U}(-\sigma,\sigma).$ Initializing $h_{0}$ as $h_{0}=\mathbf{0}$, we generate a sequence of training reservoir states $h_{0},\dots,h_{T}.$ We discard the initial $N_{spin}$ training states as an initial transient and then perform a Ridge regression (with Tikhonov regularization $\beta$) to learn a linear map $W_{uh}$ such that
$W_{uh}g(h_{i})\approx u_{i}.$ (6)
$g:N_{h}\to N_{h}$ is often taken to be the identity map or simply squaring every odd indexed entry of $h_{i}.$ We assume the latter convention, following the work of Pathak et al [pathak_model-free_2018]. Once trained the reservoir dynamics can be run autonomously as
$h_{t+1}=(1-\alpha)h_{t}+\alpha\tanh\left(W_{hh}h_{t}+W_{hu}W_{uh}g(h_{t})+\sigma_{b}\mathbf{1}\right)$ (7)
to obtain a forecast of arbitrary length. A summary of tunable hyperparameters for this architecture applied to the Lorenz system are presented in Table 19. $N_{spin}=15$ for error metrics $E_{7}$ through $E_{10}$ and $N_{spin}=100$ for all other metrics.
ESNs for KS system.
RC approaches typically rely on the latent dimension $N_{h}>>N_{u}$. However, the computational cost of the previous algorithm scales roughly quadratically with $N_{h}.$ Thus, while the above approach works well for relatively small systems, without modification it does not scale well to large states such as those encountered in PDE simulations. Pathak et al. introduced a parallel reservoir approach to address this issue by dividing a high-dimensional input into $g$ lower dimensional “chunks” [pathak_model-free_2018]. A single reservoir then accepts as input only $N_{u}/g+2L$ values, where $L$ is a locality parameter that dictates the overlap of input for two adjacent reservoirs. The output of the single reservoir is only $g$ entries of the state. Since computational cost grows linearly in the number of reservoirs, this parallel approach allows for the application of RC to higher dimensional systems. Each individual reservoir is trained exactly as for the Lorenz system; there are now just $g$ reservoirs representing different regions of the domain.
Since we introduce two new hyperparameters in the parallel setup ($L$ and $g$), when we perform our hyperparameter tuning for the KS system we fix $\alpha=1$ and $\sigma_{b}=0$, following the work of Pathak et al. The complete hyperparameter search space for the KS system is given in Table 20.
hyperparameter type min (or options) max (or none)
$\alpha$ uniform 0 1
$\sigma$ loguniform 0.0001 1.0
$\sigma_{b}$ uniform 0 2
$\rho$ uniform 0.02 1
$\beta$ loguniform $10^{-10}$ $10^{-1}$
$N_{h}$ randint 500 3000
Table 19: Hyperparameter search space for the reservoir model on the Lorenz 63 system.
hyperparameter type min (or options) max (or none)
$g$ choice {16, 32, 64, 128} $\cdot$
$\sigma$ loguniform 0.0001 1.0
$L$ randint 1 10
$\rho$ uniform 0.02 1
$\beta$ loguniform $10^{-10}$ $10^{-1}$
$N_{h}$ randint 500 3000
Table 20: Hyperparameter search space for the reservoir model on the KS system.
A.4.9 Fourier Neural Operator
Neural operators are a class of machine learning models designed to learn mappings between function spaces, in contrast to the finite-dimensional Euclidean spaces typically used in conventional neural networks. Although the inputs and outputs are discretized in practice, neural operators aim to generalize across discretizations and treat functions as the primary objects of learning.
The Fourier Neural Operator (FNO), in particular, is a neural operator architecture that replaces the kernel integral operator with a convolution operator defined in Fourier space, which allows for learning of operators in the frequency domain. It maps the input to the frequency domain using the Fourier transform, applies spectral convolution by multiplying learnable weights with the lower Fourier modes, and maps the result back to the physical domain via the inverse Fourier transform. This allows the model to learn families of PDEs, rather than solving individual instances. Without the high cost of evaluating integral operators, it maintains competitive computational efficiency.
Let $D\subset\mathbb{R}^{d}$ be a bounded domain. We consider learning an operator $G$ that maps between function spaces:
$G:\mathcal{A}\to\mathcal{U}$ (8)
where $\mathcal{A}=L^{2}(D;\mathbb{R}^{d_{a}})$ is the input function space and $\mathcal{U}=L^{2}(D;\mathbb{R}^{d_{u}})$ is the output function space.
Given an input function $a\in\mathcal{A}$, the FNO approximates the operator $G$ through a kernel integral operator:
$G(a)(x)=\sigma\left(Wa(x)+b+\int_{D}\kappa(x,y)a(y)\,dy\right)$ (9)
where $W\in\mathbb{R}^{d_{u}\times d_{a}}$ is a linear transformation, $b\in\mathbb{R}^{d_{u}}$ is a bias term, $\kappa:D\times D\to\mathbb{R}^{d_{u}\times d_{a}}$ is a learnable kernel function, and $\sigma:\mathbb{R}^{d_{u}}\to\mathbb{R}^{d_{u}}$ is a pointwise non-linear activation function.
The kernel is parameterized in Fourier space as:
$\kappa(x,y)=\sum_{k\in\mathbb{Z}^{d}}\widehat{\kappa}(k)e^{2\pi ik\cdot(x-y)}$ (10)
where $\widehat{\kappa}(k)$ are the Fourier coefficients of the kernel. The translation-invariant kernel $\kappa(x,y)=\kappa(x-y)$ enables efficient convolution. This leads to the implementation:
$G(a)(x)=\sigma\left(Wa(x)+b+\sum_{k\in\mathbb{Z}^{d}}\widehat{\kappa}(k)\widehat{a}(k)e^{2\pi ik\cdot x}\right)$ (11)
where $\widehat{a}(k)$ represent the Fourier coefficients of the input function $a$. In practice, the sum over $k\in\mathbb{Z}^{d}$ is truncated to a finite number of low-frequency modes.
Model Architecture
The architecture (Figure 8) begins with an initial fully connected multilayer perceptron (MLP) that projects the input to a higher-dimensional space, followed by four Fourier layers, and concludes with two fully connected MLPs that project the output to the desired dimensions.
Each Fourier layer performs a spectral convolution by first transforming the data into the frequency domain using Fast Fourier Transform (FFT), then multiplying the Fourier coefficients with learable weights in the frequency space, and finally transforming back to physical space using inverse FFT. The Fourier layer only keeps a limited number of the lower Fourier modes, with high modes being filtered out. Additionally, each layer adds a linearly transformed version of its input to the output of the spectral convolution, which helps preserve local features and adds flexibility to the layer’s expressiveness. Every Fourier layer is followed by a GELU activation function to introduce non-linearity.
Figure 8: Architecture of the Fourier Neural Operator from [li2021fourier]
Hyperparameters
Based on our implementation of the FNO model, which closely follows that of the original paper, we test the hyperparameters as shown in Table 21. The number of Fourier modes is tuned separately for each mode.
hyperparameter type range or options
Fourier modes integer [8,32]
Network width integer [32, 128]
Batch size choice 16, 32, 64, 128
Learning rate (lr) loguniform [0.0001, 0.01]
Table 21: Hyperparameter search space for the FNO model.
A.4.10 Kolmogorov-Arnold Networks
Kolmogorov–Arnold Networks (KANs) are a recently proposed alternative to traditional Multi-Layer Perceptrons (MLPs) [liu2024kan]. With learnable activation functions placed on edges that replace linear weights, KANs have been shown to provide improved accuracy and greater interpretability compared to traditional methods.
KANs were inspired by the Kolmogrov-Arnold representation theorem which posits that any multivariate continuous function $f$ on a bounded domain can be expressed as a finite composition and addition of univariate continuous functions [kolmogorov1957representations]. In other words, for a smooth function $f:[0,1]^{n}\to\mathbb{R}$,
$f(\mathbf{x})=f(x_{1},x_{2},...,x_{n})=\sum^{2n+1}_{q=1}\Phi_{q}\left(\sum^{n}_{p=1}\phi_{q,p}(x_{p})\right)$ (12)
where $\phi_{q,p}:[0,1]\to\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$.
Model Architecture
While the Kolmogrov-Arnold representation theorem is restricted to a small number of terms and only two hidden layers, this theorem can be generalized to increase the width and depth of the network. A single KAN layer is defined as a matrix of 1D functions thus the inner and outer functions in Equation 12, $\phi_{q,p}$ and $\Phi_{q}$, each represent a single KAN layer. A deeper network can be constructed by adding more KAN layers. A general KAN network with $L$ layers can be represented as a composition of $L$ functions:
$f(\mathbf{x})=\sum^{n_{L-1}}_{i_{L-1}=1}\phi_{L-1,i_{L},i_{L-1}}\left(\sum^{n_{L-2}}_{i_{L-2}=1}\cdots\left(\sum^{n_{2}}_{i_{2}=1}\phi_{2,i_{3},i_{2}}\left(\sum^{n_{1}}_{i_{1}=1}\phi_{1,i_{2},i_{1}}\left(\sum^{n_{0}}_{i_{0}=1}\phi_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\right)\cdots\right)$
where $n_{l}$ is the number of nodes in the $l^{th}$ layer and $\phi_{l,j,k}$ is the activation function that connects the $k^{th}$ neuron in the $l^{th}$ layer to the $j^{th}$ neuron in the $l+1$ layer. The network architecture is better illustrated in Figure 9 which was adapted from Figure 2.2 in [liu2024kan].
Figure 9: Sample architecture of a Kolmogorov-Arnold Network with three layers of size $[2,3,1]$. Activation functions $\phi$ are placed on the edges and are parametrized as a spline. Each output of a node is a sum of its inputs.
Each activation function is comprised of a basis function $b(x)$ and a spline function:
$\phi(x)=w_{b}b(x)+w_{s}\mathrm{spline}(x)$
where
$b(x)=\mathrm{silu}(x)=\frac{x}{1+e^{-x}}$
$\mathrm{spline}(x)=\sum_{i}c_{i}B_{i}(x)$
Initially, $w_{s}$ is set to $1$ and $\mathrm{spline}(x)\approx 0$. The weights of the basis function is initialized according to Xavier initializations.
KAN Implementation
Although KANs have primarily been applied to science-related tasks such as function approximation and PDE solving, Example 14 of the pykan package demonstrates their use in a supervised learning setting. In this work, the KAN implementation from that example was adapted to address the reconstruction and forecasting tasks posed in the Common Task Framework.
For forecasting tasks, the input-output pairs were constructed in an autoregressive manner, where each input consisted of lagged observations used to predict future values. The input and output dimensions depend on both the number of spatial dimensions in the dataset and the chosen lag.
The Lorenz 63 system is a three-dimensional dynamical system. For a lag of $l$, the input dimension was set to $d_{\text{in}}=3l$. While prediction windows greater than 1 were tested during training, a prediction window of 1 produced the best results. Therefore, the output dimension was fixed at $d_{\text{out}}=3$.
For the Kuramoto–Sivashinsky (KS) dataset, which contains 1024 spatial points, the input dimension was set to $d_{\text{in}}=1024l$ and the output dimension to $d_{\text{out}}=1024$.
For reconstruction tasks, the model was trained in an autoencoding fashion, where each input was mapped directly to itself as the target output. For the Lorenz 63 system, the input and output dimensions were both set to $d_{\text{in}}=d_{\text{out}}=3$. For the Kuramoto–Sivashinsky (KS) system, the dimensions were set to $d_{\text{in}}=d_{\text{out}}=1024$.
Hyperparameters
Based on the hyperparameter settings provided in the pykan package and the results reported in the original paper [liu2024kan], the hyperparameters outlined in Tables 22 and 23 were selected and tuned for this model. Broadly, the hyperparameters fall into two categories: (1) model architecture and (2) training.
Architecture-related hyperparameters include the number of layers, dimensions of hidden layers, grid resolution, the polynomial degree of the spline basis ($k$), and the lag. Training-related hyperparameters include the number of training steps (epochs), learning rate, overall regularization strength ($\lambda$), and the regularization coefficient for the spline parameters ($\lambda_{coef}$).
hyperparameter type min (or options) max (or none)
steps randint 50 $10^{4}$
$\text{lag}^{*}$ randint 1 5
lr loguniform $10^{-5}$ $10^{-1}$
num_layers randint 1 5
{one-five}_dim∗∗ randint 1 10
grid randint 1 100
k randint 1 3
$\lambda$ loguniform $10^{-7}$ $10^{-3}$
$\lambda_{coef}$ loguniform $10^{-7}$ $10^{-3}$
Table 22: Hyperparameter search space for the KAN model on the Lorenz 63 system. NOTE: The lag parameter is set to zero for reconstruction tasks (pair_id = 2 or 4). The dimension of each layer is defined separately. For example the number of nodes in layer two would be defined as two_dim∗∗.
hyperparameter type min (or options) max (or none)
steps randint 50 $10^{4}$
$\text{lag}^{*}$ randint 1 2
batch choice {-1, 50-100} $\cdot$
lr loguniform $10^{-5}$ $10^{-1}$
num_layers randint 1 5
{one-five}_dim∗∗ randint 1 10
grid randint 1 100
k randint 1 3
$\lambda$ loguniform $10^{-7}$ $10^{-3}$
$\lambda_{coef}$ loguniform $10^{-7}$ $10^{-3}$
Table 23: Hyperparameter search space for the KAN model on the KS system. NOTE: The lag parameter is set to zero for reconstruction tasks (pair_id = 2 or 4). The dimension of each layer is defined separately. For example the number of nodes in layer two would be defined as two_dim∗∗.
A.4.11 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs), introduced by Raissi et al. [raissi2019physics], have emerged as a powerful framework for solving differential equations using deep learning. Unlike standard neural networks, PINNs embed physical laws directly into the loss function, enabling them to honor both data fidelity and governing equations. The loss function is typically composed of two terms:
$\mathcal{L}(\theta,\gamma)=\mathcal{L}_{\text{data}}(\theta)+\lambda\mathcal{L}_{\text{DE}}(\theta,\gamma)=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left\lVert u_{\theta}(x_{i},t_{i})-u(x_{i},t_{i})\right\rVert_{2}^{2}+\lambda\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\left\lVert\mathcal{N_{\gamma}}[u_{\theta}(x_{i},t_{i})]\right\rVert_{2}^{2},$
Here, $u_{\theta}(x,t)$ denotes a neural network approximation of the solution with fitting parameters $\theta$, and independent variable inputs $(x,t)$. $u(x,t)$ is the ground truth at data points $(x,t)$, and $\mathcal{N}_{\gamma}[u]=0$ represents the residual, with differential operator $\mathcal{N}_{\gamma}$ and fitting model parameters $\gamma$. The first term, $\mathcal{L}\text{data}$, ensures agreement with observed data (e.g., initial and boundary conditions), while the second term, $\mathcal{L}\text{DE}$, enforces consistency with the known physical laws through collocation points.
PINNs were originally designed as differential equation solvers [lagaris1998artificial], and they excel at interpolating solutions within a domain where collocation points are defined. Their primary strength lies in approximating solutions to known equations. While they can, in principle, be extended to infer unknown parameters of the governing equations by treating them as learnable variables in the loss function, this joint optimization (i.e. over both the neural network parameters $\theta$ and the model parameters $\gamma$) is notoriously difficult. In complex spatio-temporal settings, this often leads to poor convergence and suboptimal solutions, as observed in our CTF. Recent extensions show promising directions for improvement [wang2024respecting, chen2021physics].
Implementation.
We use the DeepXDE library [deepxde] to implement the PINN architecture, building on the inverse modeling example provided for the Lorenz system [deepxde_lorenz_inverse]. In our implementation, we assume a parametric form of the target differential equation (e.g., Lorenz or Kuramoto–Sivashinsky) and treat all coefficients as learnable parameters.
Hyperparameters.
Our hyperparameter search includes the learning rate, network depth and width, and the number of training, boundary, and collocation points used to evaluate the data and physics loss terms. Table 24 summarizes the hyperparameter search space.
hyperparameter type range (or options)
Number of layers integer $[3,6]$
Number of neurons per layer integer $[10,40]$
Number of boundary points integer $[200,1000]$
Number of domain points (for PDE) integer $[200,1000]$
Learning Rate loguniform $10^{-5},10^{-2}]$
Table 24: Hyperparameter search space for PINNs.
A.4.12 Neural-ODE
Nerual-ODEs are a type of neural network that uses an ODE solver to model the hidden state of a neural network.[chen2018neural]. This is very similar to ODE-LSTMs, another model evaluated in this work, except it makes use of a vanilla MLP instead of LSTM.
We search over the following hyperparameters: hidden_state_size (dimension of the latent space), seq_length (input sequence length), batch size, and lr (learning rate).
hyperparameter type min (or options) max (or none)
hidden_state_size randint 8 1024
seq_length randint 5 74
batch_size randint 5 120
lr log_uniform $10^{-5}$ $10^{-2}$
Table 25: Hyperparameter search space for Neural-ODE models. We train for 100 epochs.
A.4.13 LLMTime
LLMTime [gruver2023large] is a time-series foundation model that uses pre-trained LLMs to perform zero-shot forecasting of time-series data. Their approach is to modify the tokenization of each model so that time-series forecasting is casted as a next-token prediction in text problem. For our evaluation, we used the llama-7b as LLMTime’s base LLM and used the default temperature of 1.0, an alpha of 0.99, and a beta of 0.3. We also used LLMTime’s default Llama tokenizer. LLMTime is only able to forecast univariate time-series, so we auto-regressively forecast each dimension with a context of 200 tokens and a prediction length of 100 tokens at a time. Once each dimension has been forecasted, they are concatenated and evaluated on the test set. For reconstruction tasks, we take the first 10 time-steps of the training data and forecast each dimension until we have a vector containing the same number of timesteps as in the testing dataset and then concatenate and calculate our metrics as before.
A.4.14 Chronos
Chronos [ansari2024chronoslearninglanguagetime] is a pre-trained probabilistic time-series foundation model from Amazon. The model is informed by the success of transformers and LLMs, and as such tokenizes time series values using scaling and quantization and trains using the cross-entropy loss function. The model is only capable of doing univariate time-series forecasting. For our evaluation, we use the pre-trained chronos-t5-base model and do a one-shot forecast of each dimension of each dataset independently and concatenate them when calculating our metrics. For reconstruction tasks, we take the first 10 time-steps of the training data and forecast each dimension until we have a vector containing the same number of timesteps as in the testing dataset and then concatenate and calculate our metrics as before. Chronos has a much smaller context length than LLMTime due to requiring more VRAM for inference.
A.4.15 Moirai
Moirai_MoE [liu2024moiraimoeempoweringtimeseries] is a time-series forecasting foundation model from Salesforce AI Research. The model uses a sparse mixture-of-experts transformer architecture and is able to do one-shot multivariate time-series forecasting on arbitrary time-series datasets. For our evaluation, we used the pre-trained base model and predicted 10 time-steps at a time with a context length of 20. For reconstruction tasks, we take the first 10 time-steps of the training data and forecast until we have a matrix containing the same number of timesteps as the testing dataset. Moirai_MoE has a much smaller context length than LLMTime due to requiring more VRAM for inference.
A.4.16 Sundial
Sundial [liu2025sundialfamilyhighlycapable] is a family of native, flexible and scalable time-series foundation models from Tsinghua University, tailored specifically for time series analysis. It is pre-trained on TimeBench (about one trillion time points), adopting a flow-matching approach rather than fixed parametric densities. Sundial directly models the distribution of next-patch values in continuous time-series without discrete tokenisation; it is built on a decoder-only Transformer architecture. For our evaluation, we used the pre-trained sundial-base-128m model; the model can handle multivariate time-series forecasting directly. For the KS evaluation, due to RAM limitations, we have split the "spatial" dimension into batches, forecasting each batch independently and concatenating the results. For reconstruction tasks, we take some of the first time-steps of the training data (around 10%) and forecast until we have a matrix containing the same number of timesteps as the testing dataset.
A.4.17 Panda
Panda [lai2025panda] is a foundation model for nonlinear dynamical systems based on Patched Attention for Nonlinear DynAmics. Panda is motivated by dynamical systems theory and adopts an encoder-only architecture with a fixed prediction horizon. It is pre-trained purely on a synthetic dataset of $2\times 10^{4}$ chaotic dynamical systems, discovered using a structured algorithm for dynamic systems discovery introduced in the same work. For our evaluation, we used the pretrained model weights provided on the official code repository associated with [lai2025panda]. The main free parameter in the forecasts with Panda is the context length. In the Lorenz evaluation we allow this to be the full dataset that we provide, but due to RAM limitations for the KS dataset we have to limit the context to 512 observations.
A.4.18 TabPFN-TS
TabPFN for Time Series (TabPFN-TS) [hoo2025tablestimetabpfnv2outperforms] is based on the tabular foundation model TabPFNv2 [hollmann_accurate_2025], adapted to the task of time series forecasting. We use the pretrained model weights, leaving the only remaining parameter as the amount of data for each specific system that the model is exposed to before performing zero-short forecasting. In the case of the Lorenz system, this is the entirety of the available training data for the task. However, for the KS system, we restrict to at most 500 time steps to be used for context. This restriction was introduced as a result of limited available memory, and is similar to the restriction placed on Panda.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper does exactly as stated in the abstract: We build a platform for evaluation scientific machine learning models on diverse challenges in science and engineering.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a separate section in the paper which clearly outlines how the CTF tasks tested are limited in scope by default as the evaluations still do not evaluate assumptions and constraints in training models. We have pointed towards how we can use this first benchmark set as a start point for future improvements.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We are benchmarking a wide range of models. The assumptions and theoretical results for each model are not applicable for this work, and well beyond the scope of what is attempted to demonstrate here: a fair comparison between methods.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes. The entire framework, datasets, and implemented methods that were scored are made available on GitHub and through Kaggle. See introduction. We implemented the ctf4science Python package to easily replicate all our results, and provide a repository with every evaluated model as a submodule that can be called from the root directory of the main repository. All configuration files used to produce the results are available in the respective model repositories and can be used to reproduce the results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Reproducibility is the core of this work. All data (https://www.kaggle.com/datasets/dynamics-ai/ctf4science-lorenz-official-ds, https://www.kaggle.com/datasets/dynamics-ai/ctf4science-kuramoto-sivashinsky-official-ds, and https://www.kaggle.com/datasets/dynamics-ai/ctf4science-sst-ds), all models, all code (https://github.com/CTF-for-Science/ctf4science) and an extensive appendix are provided to ensure full transparency, access and reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper contains all the information on the CTF. Details on the models scored on the benchmark are in the appendix, and the code to reproduce the results on their respective repositories linked above.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The merit of the CTF for science as a benchmark doesn’t depend on the statistical significance of individual scores and thus error bars were widely omitted. Despite this, our main result in Table 1 and Figure 3 provide the mean and standard deviations over five full training then evaluation runs on the test set. We also include error bars in Fig. 3 and 4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See section 1.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We ensured full compliance.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the impact on the scientific community, but not society as a whole. We consider this work benign in nature and thus focused our discussion on the groups of people directly affected by ctf4science in the near term: researchers, academics, and engineers.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: We consider the datasets and framework of ctf4science benign and don’t see high-risk for misuse or dual use at this time.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All sources and assets were cited appropriately.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Datasets are documented in the paper and provided in the croissant format. Code is documented and made publicly available. Modeling methods used and implemented are documented extensively in the appendix and in their respective repositories linked above.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Not applicable
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Not applicable
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Not applicable
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.